### v2

# A Text-guided Protein Design Framework

Abstract
Current AI-assisted protein design mainly utilizes protein sequential and structural information. Meanwhile, there exists tremendous knowledge curated by humans in the text format describing proteins’ high-level functionalities. Yet, whether the incorporation of such text data can help protein design tasks has not been explored. To bridge this gap, we propose ProteinDT, a multi-modal framework that leverages textual descriptions for protein design. ProteinDT consists of three subsequent steps: ProteinCLAP which aligns the representation of two modalities, a facilitator that generates the protein representation from the text modality, and a decoder that creates the protein sequences from the representation. To train ProteinDT, we construct a large dataset, SwissProtCLAP, with 441K text and protein pairs. We quantitatively verify the effectiveness of ProteinDT on three challenging tasks: (1) over 90% accuracy for text-guided protein generation; (2) best hit ratio on 10 zero-shot text-guided protein editing tasks; (3) superior performance on four out of six protein property prediction benchmarks.
Machine learning (ML) has recently shown profound potential for protein discovery. These ML tools have been quickly adapted as auxiliary and accelerating roles in scientific pipelines, including but not limited to protein engineering [1], structure prediction [2], structure reconstruction [3], and inverse folding [122].
To this end, most existing such methods merely model the intrinsic properties of protein sequences and structures [101, 6, 102, 8, 9, 10, 11]. Motivated by recent breakthroughs of foundation models [86, 88, 89, 90], approaches in the computational chemistry domain that utilize textual descriptions of drugs in addition to their intrinsic chemical and structural information have proved to be effective in small-molecule drug discovery [16, 95, 96, 97, 138]. This naturally raises a question for the protein domain: how can such a multi-modal mechanism benefit protein engineering and generation? To answer this question, we initiate a novel paradigm that leverages textual descriptions for protein design.
Specifically, we aim to exploit the following two modalities of proteins: the protein sequence and the textual description. The protein sequence is the sequence of 20 types of amino acids (a.k.a. residues) that influence how the protein will fold and what functions it can perform; while the second modality is the textual description recorded in the open data sources, e.g., UniProt [21]. Such annotation data can contain abundant knowledge of proteins [22], such as biological processes they participate in, molecular functions they execute, and cellular components they localize to. These two modalities respectively focus on expressing the internal biochemical composition and the high-level knowledge summarized by domain experts. Thus, it is worth exploring the incorporation of both to realize more challenging and promising protein design tasks, such as zero-shot generalization.
Our Approach. To attain the aforementioned goal, we propose Protein Design with Text (ProteinDT), a multi-modal framework for protein design. ProteinDT is composed of three key steps, as shown in Figure 1. (1) A Contrastive LAnguage and Protein pretraining (ProteinCLAP) step to align the representations between text sequences and protein sequences. ProteinCLAP adopts the contrastive learning paradigm to align the representation space of the two modalities. This is enabled by a constructed dataset called SwissProtCLAP, with around 441K text-protein pairs extracted from the SwissProt subset of UniProt. (2) A ProteinFacilitator model that produces the protein sequence representation from the text modality. We leverage a Gaussian distribution to estimate this conditional distribution. (3) A decoder model for protein generation. It is a conditional generative model that generates protein sequences conditioned on protein information, i.e., the representation produced by the previous step. Concretely, we present a robust solution to the demanding task by introducing the most powerful generative models for sequence data: one autoregressive (AR) and two diffusion models (ProteinDiff).
We consider three downstream tasks to verify the versatile functionalities of ProteinDT. The first task is text-to-protein generation, where we follow the ProteinDT pipeline to generate protein sequences from the input text prompts, which describe the desired protein attributes. Using AR and ProteinDiff, the optimal retrieval accuracy can reach over 90%. The second task is zero-shot text-guided protein editing, where the inputs are text prompts and protein sequences. We introduce two editing methods: latent interpolation and latent optimization. Latent interpolation conducts an interpolation in the sequence-level representation space, while latent optimization directly optimizes the token-level representation. Both editing methods inject the text modality and the learned representation is used for protein generation accordingly. We observe that ProteinDT can reach consistently the best hit ratio on 10 editing tasks, ranging from structure- and stability-aware tasks, to peptide-binding editing tasks. Additional qualitative results offer further compelling evidence supporting the validity of ProteinDT. The last task is protein property prediction, aiming to evaluate the robustness and generalization ability of the learned representation from ProteinCLAP. Compared to six state-of-the-art protein sequence representation methods, ProteinDT can obtain consistently superior performance on four of six benchmark tasks.
Figure 1: Pipeline for ProteinDT framework (a-c) and downstream tasks (d-f). (a) ProteinCLAP, a contrastive learning paradigm, aligns the representation space of the text and protein sequence modalities. (b) ProteinFacilitator model augments the mapping from text sequence representation to protein sequence representation. (c) A protein sequence decoder, which generates protein sequences conditioned on the representations produced from previous steps. (d) Downstream text-to-protein generation task. (e) Downstream text-guided protein design task. (f) Downstream protein property prediction task.
Results
In this section, we first provide the preliminaries of ProteinDT and dataset construction. Then we assess three downstream tasks to verify the versatile functionalities of ProteinDT. We focus on exploring two attributes of ProteinDT: (1) ProteinDT can enable zero-shot protein design and editing based on textual descriptions, and (2) ProteinDT can learn a robust protein representation. The specific descriptions for the three downstream tasks, model implementation details, and training hyperparameters can be found in Supplementary B and C.
Dataset Construction and Two Modalities
We construct SwissProtCLAP, a text-protein pair dataset, from UniProt [21]. UniProt is the most comprehensive knowledge base for protein sequences. It also contains extensive annotations including function descriptions, domains, subcellular localization, post-translational modifications, and functionally characterized variants [23]. SwissProt [24], originally created in 1986, is now the subset of UniProt that contains manually reviewed information and expert curation.
Based on SwissProt, we collect and curate a dataset called SwissProtCLAP, with 441,013 text-protein sequence pairs in total. The 441K proteins can be categorized into 327,577 genes and 13,339 organisms. One concrete example can be found in Figure 1, and more details are in Supplementary B.
We also note that the dataset size of SwissProtCLAP is relatively small compared to the peers from other domains (e.g., 400M text-image pairs in the language-vision domain [86]). To alleviate the data insufficiency issue, instead of training our model from scratch, we adopt pretrained representation encoders to train ProteinDT, as elaborated below.
Protein sequence ${\bm{x}}_{\text{p}}$ and representation ${\bm{z}}_{\text{p}}$. Proteins have four levels of structure [25]: primary, secondary, tertiary, and quaternary structure, and the corresponding degree of complexity in the polypeptide chain increases. In this work, we are interested in the protein primary structure ${\bm{x}}_{\text{p}}$, i.e., the amino acid sequence. There are in total 20 different amino acids in proteins, represented with 20 characters, as shown in Figure 1. For the protein sequence representation ${\bm{z}}_{\text{p}}=f_{\text{p}}({\bm{x}}_{\text{p}})$, we use BERT as the encoder [26]. It is one of the most advanced Transformer models for modeling sequential data [27]. Additionally, we take the BERT model with ProtBERT pretraining [102]: it was pretrained on a large protein sequence corpus [28, 29], and the pretraining task was to reconstruct the masked tokens.
Text sequence ${\bm{x}}_{\text{t}}$ and representation ${\bm{z}}_{\text{t}}$. The textual descriptions ${\bm{x}}_{\text{t}}$ are formulated in the form of sequences of tokens, and we also adopt the BERT model as the encoder ${\bm{z}}_{\text{t}}=f_{\text{t}}({\bm{x}}_{\text{t}})$. We further adapt the pretrained SciBERT [132], which was pretrained on the full-text computer science and biomedical papers from Semantic Scholar [31]. Notice that although both modalities use BERT-like models, the detailed architectures and token vocabulary are different. Further, such a disentangled design enables us to adopt the respective pretrained models (e.g., ProtBERT and SciBERT) to mitigate the limited availability of structured text-protein sequence pairs for training.
Pretraining: ProteinDT
The core idea of ProteinDT is the integration of textual descriptions and protein sequences. Such jointly learned knowledge empowers ML models to form informative protein representations with perception for both modalities. As a result, these representations not only contain text-centric domain knowledge about proteins, but also enable the tuning of the language descriptions to accomplish the text-guided protein generation and optimization tasks.
As illustrated in Figure 1, through leveraging the constructed SwissProtCLAP dataset, ProteinDT attains the aforementioned goal with three key steps: a pretraining model, a facilitator, and a decoder. First, the ProteinCLAP model in ProteinDT utilizes contrastive learning to construct representations that integrate knowledge from both text and protein sequences. By doing so, ProteinCLAP not only forms an informative representation of proteins, but also enables the text-protein representation alignment, enabling rich cross-modality tasks like text-protein retrieval. The second step is to train a ProteinFacilitator model, which is an augmented alignment module in addition to ProteinCLAP. To be more concrete, the ProteinFacilitator model maps a piece of text prompt (e.g., “A protein that is stable.") to a corresponding protein representation that captures both the semantic information of the given text and essential protein sequence patterns. Last but not least, ProteinDT subsequently utilizes a conditional generative model to decode the formed protein representation into the protein sequence. We propose one autoregressive model (AR) with Transformer and two diffusion models (ProteinDiff). More details of three pretraining steps can be found in the Methods section. With such a ProteinDT framework, we can explore a broad range of challenging protein design tasks, as will be illustrated below.
Downstream Task: Text-to-Protein Generation
To illustrate that natural language can guide protein semantics through ProteinDT, we propose the text-to-protein generation task. The pipeline is shown in Figure 1. For the decoder, we consider a Transformer-based AR model and two ProteinDiff models (RNN and Transformer as the transition network, respectively). Notice that we have two options for the condition in decoding. (1) Without ProteinFacilitator: For the input text sequence, we initially obtain the text sequence representation using ProteinCLAP. Subsequently, we directly utilize this representation as a condition to generate the protein sequence using the pretrained encoder. (2) With ProteinFacilitator: Alternatively we encode the input text sequence and employ ProteinFacilitator to derive the protein representation. This modified protein representation is then employed as a condition for generating the protein sequence. An illustration is in Figure 1.
Figure 2: Visualization of text-to-protein generation and text-guided protein editing. (a) Visualization for evaluation of text-to-protein generation. The pretrained ProteinCLAP is used to calculate the similarity between the sampled text and generated protein sequence pairs. (b-c) Two methods for text-guided protein editing: latent interpolation and latent optimization. (d-f) Visualization for evaluation of text-guided protein editing. For three types of editing tasks, different evaluation metrics (marked in red) are applied accordingly.
Table 1: Retrieval accuracy (%) for text-to-protein generation. It is a multiple-choice task ($T$ options), measuring the consistency between the input text prompt and the generated protein sequence.
Galactica ChatGPT ProteinDT (Ours) w/o ProteinFacilitator w/ ProteinFacilitator AR AR AR ProteinDiff AR ProteinDiff Transformer RNN Transformer Transformer RNN Transformer T = 4 51.5 38.5 49.00 24.00 35.50 97.00 40.50 51.50 T = 10 29.0 23.0 27.00 10.50 17.50 91.00 21.50 25.00 T = 20 19.0 15.5 20.00 5.50 9.50 83.50 15.00 13.50
Experiment setting. We randomly sample 200 text sequences from SwissProtCLAP as the text prompts and then generate a protein sequence for each following the ProteinCLAP-ProteinFacilitator-decoder pipeline and ProteinCLAP-decoder pipeline. For evaluation, we adopt the well-trained ProteinCLAP to obtain the retrieval accuracy: each protein sequence is compared to the input text prompt and $T-1$ randomly chosen texts using ProteinCLAP, and a match is successful if the actual text-sequence pair ranks the highest among all comparisons. The evaluation pipeline is shown in Figure 2.
Baselines. We consider two large language models as baselines. Galactica (1.3b) [32] is specifically pretrained for scientific tasks. ChatGPT (3.5-Turbo-0301) is one of the most general-purpose foundation models, and existing works have shown its potential in drug editing tasks [138, 33, 34].
Observations. The main results are presented in Table 1, revealing three key observations. Observation 1. Adopting the ProteinFacilitator model is deemed essential. As shown in Table 1, across all ProteinDT decoder models, incorporating ProteinFacilitator yields higher accuracies compared to models without it. Especially for decoders like AR-Transformer, the performance gaps are substantial by up to around 50%. Observation 2. For the three decoder variants of ProteinDT, AR-Transformer can beat two diffusion models (ProteinDiff). Similar findings are also reported in recent studies on text-to-text generation [35, 36]. We speculate that either the AR model excels in fitting discrete and sequential data models, or we haven’t found the most expressive diffusion model for such a challenging task. These are interesting ML research questions for future exploration. Observation 3. When compared to baseline models, the AR-Transformer with ProteinFacilitator reaches consistently better performance, while the two variants of ProteinDiff with ProteinFacilitator obtain similar performance than Galactica and ChatGPT. It is noteworthy that both baselines are Transformer-based autoregressive models.
Downstream Task: Zero-shot Text-guided Protein Editing
The next task we consider is zero-shot text-guided protein editing. The inputs are an initial protein sequence and a text prompt describing the target property, and the output is the edited protein sequence. The textual modality enables ProteinDT to adapt to flexible target tasks for editing in a zero-shot manner, as long as the text prompt is available. Next, we introduce two text-guided protein editing methods: latent interpolation and latent optimization.
Latent interpolation. We have two representations for the text prompt and input protein as ${\bm{z}}_{\text{t}}$ and ${\bm{z}}_{\text{p}}$, respectively. Then we apply the spherical linear interpolation (slerp) [89], which yields an intermediate representation $\tilde{\bm{z}}_{\text{c}}=\text{slerp}({\bm{z}}_{\text{t}},{\bm{z}}_{\text{p}},\theta)$, and the output protein is generated from decoder, as $\tilde{\bm{x}}_{\text{p}}\sim p({\bm{x}}_{\text{p}}|\tilde{\bm{z}}_{\text{c}})$. Here $\theta\in[0,1]$ is a coefficient controlling how close $\tilde{\bm{x}}_{\text{p}}$ is to the text prompt or the input protein, and the larger value of $\theta$ expects $\tilde{\bm{x}}_{\text{p}}$ to be closer to the text prompt.
Latent optimization. We also propose using latent optimization [90, 97] for text-guided protein editing. The high-level idea is to learn a token-level latent code ${\bm{z}}_{w}$ that is close to both the text and protein sequence-level representation, as ${\bm{z}}_{w}=\arg\min(\lambda\|P({\bm{z}}_{w})-{\bm{z}}_{\text{t}}\|^{2}+(1-\lambda)\|P({\bm{z}}_{w})-{\bm{z}}_{\text{p}}\|^{2})$, where $P$ is the pooling function, and $\lambda$ controls the closeness to the target prompt. Then, the learned ${\bm{x}}_{w}$ are mapped to the protein sequence using a pretrained decoder $g$, i.e., $\tilde{\bm{x}}_{\text{p}}=g({\bm{z}}_{w})$. Notice that $g$ is a token-wise decoder, and can be trained in an auto-encoder manner with the fixed ProteinCLAP protein encoder. More details of the two methods are in Supplementary C.
Editing tasks and evaluation metrics. To fully verify the effectiveness of ProteinDT, we design 10 zero-shot text-guided protein editing tasks. We further categorize them into three settings as detailed below. For editing evaluation, we employ the satisfactory hit ratio (%), which measures the ratio of the output score being higher or lower than the input score, as indicated in the text prompt. The scores are provided by oracle evaluators on each protein sequence, as discussed below. To mitigate bias resulting from different oracle evaluators, we adopt two distinct types of evaluators for each task. Certain evaluations require proteins’ 3D structures, for which we employ the ESMFold algorithm [8]. The pipeline of the task design and evaluation are in Figure 2.
(1) Structure editing. Our initial focus involves four editing tasks associated with the secondary structure [138]. The prompts are in the format of “modify the amino acid sequence to have more (less) alpha helices” and “modify the amino acid sequence to have more (less) beta sheets”. We adopt two methods to measure the secondary structures for each amino acid: (1) A strong protein sequence evaluator for secondary structure prediction (as will be discussed in the property prediction downstream below); (2) Define Secondary Structure of Proteins (DSSP) algorithm for assigning secondary structure to the amino acids [37].
(2) Stability editing. Then we consider four types of stability editing tasks. Two proteins are considered [137]: villin HP35 and Pin1 WW domain. The prompts are in the format of “modify the amino acid sequence to have higher (lower) stability”. We consider two evaluators: (1) A strong protein sequence evaluator for stability (as will be discussed in the property prediction downstream below); (2) The predicted local distance difference test (pLDDT) score measuring the confidence for folding, which has been used as a proxy for stability [2].
(3) Peptide binding editing. The last two tasks are peptide binding editing. Peptides are small proteins (usually less than 15 amino acids) with specific functionalities through binding to the target proteins. Here we take 154 pairs of peptide-protein binding pairs from Protein Data Bank (PDB) [39], where the proteins have textual descriptions from UniProt in SwissProtCLAP. Then, the task is to make the input peptide with higher/lower binding affinity with respect to the target protein (textual description). For scoring with the oracle evaluator, we consider two options: (1) A pretrained deep learning model for binding affinity prediction [140]; (2) GNINA to calculate the docking score before and after editing [142].
Baselines. As in the text-to-protein generation tasks, we use Galactica (1.3b) and ChatGPT (3.5-Turbo-0301) for editing baselines. The prompts for ProteinDT and baselines may be modified accordingly. Please refer to Supplementary D for detailed prompts.
Observations. The main results are in Table 2. We can observe that two variants of ProteinDT (latent interpolation and latent optimization) generally perform better on all 10 tasks and 20 evaluation methods. However, for the stability editing tasks, latent interpolation methods tend to favor lower stability, while latent optimization can reach a good balance between higher and lower stability tasks. This suggests that latent optimization is in general a preferable method for protein editing. Despite being a generic foundation model, ChatGPT is occasionally the second-best performer on some of the stability editing tasks.
We further provide qualitative results in Figure 3. Figure 3 (a) and (d) show the structures of two input protein sequences. Then ProteinDT adopts the prompts with key words “more $\alpha$-sheets” and “more $\beta$-strands”, the edited protein structures (edited sequences folded with ESMFold) are shown in Figure 3 (b) and (e). Similarly for Figure 3 (c) and (f) when using prompts like “less $\alpha$-sheets” and “less $\beta$-strands”. For peptide editing illustration, we take PDB 3IQI, Haemophilus influenzae O-acetylserine sulfhydrylase in complex with the peptide MNENI, as an example. The input peptide is shown in Figure 3 (g), and the output peptides with higher and lower binding affinity are shown in Figure 3 (h, i). We marked the hydrogen bonds in dashed yellow lines, and there are 6, 7, and 2 bonds in Figure 3 (g, h, i), respectively.
Table 2: The satisfactory hit ratio (%) of text-guided protein editing tasks. The best results and the second-best results are marked.
Structure: $\alpha$-helices Structure: $\beta$-sheets Stability: Villin Stability: Pin1 binding editing more less more less more stable less stable more stable less stable higher lower BERT DSSP BERT DSSP BERT DSSP BERT DSSP BERT pLDDT BERT pLDDT BERT pLDDT BERT pLDDT BindingNet GNINA BindingNet GNINA Galactica 5.46 7.02 13.26 10.53 1.95 2.14 11.50 11.31 7.81 7.81 73.44 67.19 1.41 2.82 8.45 7.04 26.61 21.77 33.87 25.00 ChatGPT 5.46 6.82 7.80 6.43 4.68 2.92 5.07 3.51 29.69 17.19 78.12 18.75 28.17 18.31 39.44 36.62 16.94 12.90 26.61 35.48 ProteinDT (Ours) Latent Interpolation ProteinDiff-RNN 67.64 75.63 19.49 10.14 27.88 11.31 53.61 69.98 17.19 1.56 82.81 98.44 4.23 4.23 95.77 94.37 58.87 29.84 37.90 70.16 ProteinDiff-BERT 38.79 51.27 48.15 34.70 39.38 15.79 42.30 65.50 9.38 1.56 93.75 98.44 2.82 5.63 97.18 94.37 50.81 29.03 43.55 67.74 AR-BERT 27.49 45.61 56.92 39.38 9.16 5.26 69.79 73.10 6.25 1.56 93.75 98.44 1.41 7.04 98.59 92.96 53.23 29.84 45.97 71.77 Latent Optimization 51.66 54.19 33.92 32.94 46.59 39.57 39.38 46.59 50.00 29.69 54.69 76.56 52.11 42.25 47.89 56.34 39.52 37.90 60.48 61.29
Figure 3: Visual analysis on text-guided protein editing with latent optimization. (a-c) visualize structure editing with more/less $\alpha$-helices editing. (d-f) visualize structure editing with more/less $\beta$-sheets editing. (g-i) visualize peptide binding editing on PDB 3IQI.
Downstream Task: Protein Property Prediction
Last but not least, we assess the intermediate protein representation learned from the contrastive pretraining step using the ProteinCLAP model. ProteinCLAP aligns the text-protein sequence pairs, and the protein modality uses an encoder initialized from ProtBERT [102]. ProtBERT was pretrained on BFD in a masked language model (MLM) manner [28, 29], a self-supervised learning method to predict the masked tokens, and only the protein sequence information is included. We compare the original pretrained ProtBERT with the updated representation from ProteinCLAP.
Experiment setting. We consider the following six tasks gathered from TAPE [139]. Secondary structure (SS) prediction [43] is a token-level prediction task, and the goal is to predict the secondary structure (helix, strand, etc.) for each amino acid. SS-Q3 and SS-Q8 refer to the three-way and eight-way classifications, respectively. Homology prediction [44, 45] is to predict the protein fold of the protein sequence. The evaluation metric is accuracy for these three tasks. Contact [46, 47, 48] is to predict if pairs of amino acids within a protein sequence pair are in contact ($<8\accentset{\circ}{A}$ apart), and the metric is the precision for the top $L/2$ contacts, where $L$ is the length of the medium- and long-range proteins. Fluorescence [49] is to predict the log-fluorescence intensity of a protein. Stability [137] is to predict a proxy for protein stability. Both Fluorescence and Stability are regression tasks, and they are measured using Spearman’s $\rho$ metric. More details of the datasets are in Supplementary B.
Baselines. LSTM [50] and ResNet [51] are the two classic representation methods for sequential data. TAPE Transformer [139] is a Transformer specific for protein sequence representation. MSA Transformer [6] adds the multiple sequence alignments (MSAs) into the Transformer. The most comparable models to ProteinDT are ProtBERT [102] and its variant because ProteinCLAP shares the same backbone model with ProtBERT. OntoProtein [99] is a variant of ProtBERT and should be highlighted because (1) it also starts the pretraining from a pretrained ProtBERT, and (2) it utilizes the same modalities for pretraining, i.e., textual description and protein sequence. Yet the difference is that OntoProtein conducts a contrastive pretraining based on a knowledge graph (KG). The KG covers 565K proteins, 47K Gene Ontology (GO) terms [22], and 31 relations.
Observations. The property prediction results are in Table 3. We show two ProtBERT models followed with ProteinCLAP pretraining, corresponding to the EBM-NCE loss and InfoNCE loss, respectively. We can observe that both ProteinCLAP pretrained representations can consistently reach the best performance on three structure-related prediction tasks. Yet for the other two tasks, Homology and Stability, they are still within the top-three models. The only exception is the Fluorescence task, where all the ProtBERT models are performing worse than other baselines. Noticeably, ProteinCLAP outperforms OntoProtein on five out of six tasks.
Table 3: Results on six protein property prediction tasks from TAPE. The best results and the second-best results are marked. Baseline results are replicated from OntoProtein, with the exception of ProtBERT and OntoProtein for Contact prediction, both of which were rerun.
Method Structure Evolutionary Engineering SS-Q3 $\uparrow$ SS-Q8 $\uparrow$ Contact $\uparrow$ Homology $\uparrow$ Fluorescence $\uparrow$ Stability $\uparrow$ LSTM 0.75 0.59 0.26 0.26 0.67 0.69 TAPE Transformer 0.73 0.59 0.25 0.21 0.68 0.73 ResNet 0.75 0.58 0.25 0.17 0.21 0.73 MSA Transformer - 0.73 0.49 - - ProtBERT 0.81 0.67 0.59 0.29 0.61 0.82 OntoProtein 0.82 0.68 0.56 0.24 0.66 0.75 ProteinDT-ProteinCLAP-InfoNCE (Ours) 0.8331 0.6935 0.6031 0.3092 0.6029 0.8130 ProteinDT-ProteinCLAP-EBM-NCE (Ours) 0.8334 0.6919 0.6007 0.2841 0.6105 0.7952
Discussion
We acknowledge two key limitations that may hinder ProteinDT for broader usage, as discussed below.
Challenge regarding dataset quality. In this work, we construct a text-protein pair dataset called ProteinDT. It has 441K data pairs, yet such a dataset size is small compared to the vision and language domains. If we want to take the protein backbone structure into consideration, that would be reduced to merely 45K pairs of data [100]. Thus, data insufficiency has become the bottleneck of this research direction. ProteinDT solves this by adopting pretrained checkpoints, while more promising solutions exist. For example, it may be possible to create alternative datasets from full-text biomedical manuscripts by using named entity recognition to identify proteins and mapping their names to identifiers and sequences [54]. This approach may generate a larger, more diverse, but lower-quality text corpus paired with a more restricted set of proteins.
Challenge regarding evaluation. For the text-guided protein design tasks, the textual description is quite flexible, but the output proteins can be tested. Although most protein properties can be tested experimentally, the cost and throughput of the experiments can vary considerably depending on the property. Generative models in the art, language, and vision domains are often subjective and can be evaluated by a human reviewer. In contrast, protein engineering is objective but requires wet lab testing for the most rigorous model assessment. In this work, we take the computationally feasible solutions with regressor models as surrogate oracles, following the evaluation approach of reinforcement learning models [114]. However, the validity of such surrogate models is not guaranteed. We attempt to reduce our dependence on surrogate models during the editing evaluation by including alternative evaluations from methods such as ESMFold and GNINA, but those tools still do not provide biological ground truth. Although the preliminary protein editing results are quite encouraging, additional evaluations will be required to more thoroughly assess the strengths and weaknesses of ProteinDT on diverse editing tasks.
Conclusion. In this work, we introduced a text-guided protein design framework named ProteinDT. ProteinDT trains a ProteinCLAP model, a ProteinFacilitator model, and a decoder model (AR and ProteinDiff), to bridge the gap between the natural language and protein sequence structure. We examined and developed three downstream tasks to verify the versatile functionalities of ProteinDT: text-to-protein generation, zero-shot text-guided protein editing, and protein property prediction.
Societal impacts. Although text-based protein design has many potential positive applications in agriculture, bioengineering, and therapeutics, it can be considered a dual-use technology. Much like generative models for small molecules [56], ProteinDT could be applied to generate toxic or otherwise harmful protein sequences. Even though acting on these designs would require a wet lab, synthesizing custom amino acid sequences is typically straightforward compared to synthesizing novel small molecules. These potential negative impacts in the current version of ProteinDT may be limited by the text in the SwissProtCLAP dataset because the model is not expected to perform well for terms and concepts not represented in UniProt. Our initial manual review of a sample of SwissProtCLAP indicated it does contain phrases describing toxic proteins (e.g., “May have toxic activity. Expressed by the venom gland.”) but other potentially harmful terms examined (e.g., “dangerous”) were used in different contexts that do not pose a risk to human health (e.g., “when trehalose levels in the cell become dangerously low, TreS could expedite the conversion of glycogen to maltose via its amylase activity and then convert the maltose to trehalose”). Future efforts to expand the training dataset and model improvements could increase the dual-use risk.
Outlook. One ambitious vision of ProteinDT is to enable zero-shot protein engineering for specific protein functions. Deep learning models can guide protein design when trained on large high-throughput experimental datasets [57, 58], which are only available for a few protein functions. Specialized approaches reduce the data requirements but still require some function-specific experimental data [59, 9]. General protein language models can have a good correlation with specific functional assays in a zero-shot setting but cannot distinguish different properties of the same protein [60, 8]. ProteinDT shows a potential strategy for jump-starting future protein design of specific functions without requiring experimental data by leveraging the rich information in the textual description.
Method
Backgrounds on Generative Models
Autoregressive Model
The autoregressive (AR) model is one of the state-of-the-art generative paradigms for estimating the distribution of discrete sequential data [61, 62, 63], including text sequences and protein sequences. Suppose the data can be formulated as a sequence of $N$ components as ${\bm{x}}=({\bm{x}}^{(1)},{\bm{x}}^{(2)},\dots,{\bm{x}}^{(N)})$, such as amino acids in a protein sequence. The corresponding data distribution is factorized in an autoregressive manner: $p({\bm{x}})=p({\bm{x}}^{(1)}|{\bm{x}}^{(0)})p({\bm{x}}^{(2)}|{\bm{x}}^{(0:1)})\dots p({\bm{x}}^{(N)}|{\bm{x}}^{(0:N-1)})$, and we set ${\bm{x}}^{(0)}=\emptyset$ and $p({\bm{x}}^{(1)}|{\bm{x}}^{(0)})=p({\bm{x}}^{(1)})$ as the starting points. Parameterized by $p_{\theta}({\bm{x}}^{(n)}|{\bm{x}}^{(0:n-1)}),n\in\{1,2,\dots N\}$, the training objective is next-token prediction:
$\mathcal{L}=\mathbb{E}_{{\bm{x}}}\big{[}-\sum_{n=1}^{N}\log p({\bm{x}}^{(n)}|{\bm{x}}^{(0:n-1)})\big{]}.$ (1)
During the inference stage, running the AR model repeatedly can generate the sequence token-by-token.
Denoising Diffusion Generative Model
The key idea for the diffusion generative model is to gradually map the data to random noise, and generate samples by the parameterized denoising process. In specific, given a data point (e.g., protein sequence) from the data distribution ${\bm{x}}_{0}\sim q({\bm{x}}_{0})$, the forward process generates a Markov chain of latent variables ${\bm{x}}_{1},{\bm{x}}_{2},\ldots,{\bm{x}}_{T}$ via the following conditional distributions:
$q({\bm{x}}_{t}|{\bm{x}}_{t-1})=\mathcal{N}({\bm{x}}_{t};\sqrt{1-\beta_{t}}{\bm{x}}_{t-1},(\beta_{t})I),$ (2)
where $\{\beta_{t}\}_{t=1}^{T}$ is the variance schedule, and it keeps non-decreasing as $t\to T$. Such a forward process can be viewed as progressively adding noise to ground-truth data ${\bm{x}}_{0}$, and finally nearly leads to a complete noise ${\bm{x}}_{T}\sim p(x_{T})$.
The reverse process aims to gradually denoise a randomly sampled noise ${\bm{x}}_{T}$, so as to generate the given data point ${\bm{x}}_{0}$. To attain this goal,  [64] derives a simplified surrogate objective function, as follows:
$\mathcal{L}=\mathbb{E}_{t\sim[1,T],{\bm{x}}_{0}\sim p({\bm{x}}_{0})}\big{[}\|\epsilon_{\theta}({\bm{x}}_{t},t)-\epsilon\|^{2}\big{]},$ (3)
where $\epsilon\sim\mathcal{N}(0,I)$ is drawn from a Gaussian noise distribution and $\epsilon_{\theta}({\bm{x}}_{t},t)$ is a parameterized neural network learning to estimate the noise. The above equation is equivalent to the denoising score matching method [65, 66], and both can be united under a stochastic differential equation (SDE) framework [67]. In this work, we are interested in the discrete version of the diffusion model and we implement ProteinDT under the SDE framework to benefit future exploration.
For notation clarity, in what follows, we use ${\bm{x}}_{\text{t}}$ to represent the protein sequence data and $\tilde{\bm{x}}_{t}$ to represent the diffused data at timestep $t$ in the diffusion model.
ProteinCLAP Model
Maximizing the mutual information (MI) between modalities has been widely used for representation pretraining [68, 69]. The intuition is that MI measures the non-linear dependency between random variables (e.g., text and protein), and thus maximizing MI between modalities can force them to share more critical information. Along this research line, contrastive learning has proven its effectiveness in various tasks [70, 71, 86, 72], and the high-level idea is to transform the density estimation problem into a classification problem. Thus, we adopt it to align the representation of two modalities, coined as Contrastive LAnguage and Protein pretraining (ProteinCLAP).
To train ProteinCLAP, first, we extract positive and negative pairs from SwissProtCLAP: a text-protein pair is labeled as positive if and only if it is extracted from the same protein record in SwissProtCLAP. Then, ProteinCLAP aligns the positive pairs by increasing their similarity scores and contrasts the negative pairs by decreasing their similarity scores. In ProteinCLAP, the similarity score (e.g., cosine similarity) can be easily calculated by mapping the text and protein sequence to the representation space. The last stage is to conduct the alignment and contrasting simultaneously, where we consider two commonly-used loss functions, EBM-NCE [72] and InfoNCE [70]:
$\mathcal{L}_{\text{EBM-NCE}}=-\frac{1}{2}\mathbb{E}_{{\bm{x}}_{\text{p}},{\bm{x}}_{\text{t}},{\bm{x}}_{\text{t}}^{\prime}}\big{[}\log\sigma(E({\bm{x}}_{\text{p}},{\bm{x}}_{\text{t}})+\log(1-\sigma(E({\bm{x}}_{\text{p}},{\bm{x}}_{\text{t}}^{\prime}))\big{]}-\frac{1}{2}\mathbb{E}_{{\bm{x}}_{\text{p}},{\bm{x}}_{\text{p}}^{\prime},{\bm{x}}_{\text{t}}}\big{[}\log\sigma(E({\bm{x}}_{\text{p}},{\bm{x}}_{\text{t}})+\log(1-\sigma(E({\bm{x}}_{\text{p}}^{\prime},{\bm{x}}_{\text{t}}))\big{]},$ (4)
$\mathcal{L}_{\text{InfoNCE}}=-\frac{1}{2}\mathbb{E}_{{\bm{x}}_{\text{p}},{\bm{x}}_{\text{t}}}\Big{[}\log\frac{\exp(E({\bm{x}}_{\text{p}},{\bm{x}}_{\text{t}}))}{\exp(E({\bm{x}}_{\text{p}},{\bm{x}}_{\text{t}}))+\sum_{{\bm{x}}_{\text{t}}^{\prime}}\exp(E({\bm{x}}_{\text{p}},{\bm{x}}_{\text{t}}^{\prime}))}\Big{]}-\frac{1}{2}\mathbb{E}_{{\bm{x}}_{\text{p}},{\bm{x}}_{\text{t}}}\Big{[}\log\frac{\exp(E({\bm{x}}_{\text{p}},{\bm{x}}_{\text{t}}))}{\exp(E({\bm{x}}_{\text{p}},{\bm{x}}_{\text{t}}))+\sum_{{\bm{x}}_{\text{p}}^{\prime}}\exp(E({\bm{x}}_{\text{p}}^{\prime},{\bm{x}}_{\text{t}}))}\Big{]},$ (5)
where ${\bm{x}}_{\text{t}}^{\prime}$ and ${\bm{x}}_{\text{p}}^{\prime}$ are negative samples randomly sampled from the empirical data distribution, $E(\cdot,\cdot)$ is an energy (scalar) function defined over two modalities and we take the dot product of the representation accordingly [73], i.e., $E({\bm{x}}_{\text{t}},{\bm{x}}_{\text{p}})=\langle{\bm{z}}_{\text{t}},{\bm{z}}_{\text{p}}\rangle$. The difference between EBM-NCE (Equation 4) and InfoNCE (Equation 5) is that the former performs binary classification and the latter conducts multiclass classification. Although it has been shown that EBM-NCE is superior on the graph-related contrastive pretraining tasks [74, 72, 75], here we consider both of them for a comprehensive hyperparameter search progress.
By doing so, ProteinCLAP forms informative representations for proteins, and such robust representations help improve representation-oriented downstream tasks such as protein property prediction, as discussed in the Results Section. Another benefit of ProteinCLAP is its mapping of representations between textual descriptions and protein sequences. As a result, such a multi-modal bridge empowers retrieving protein representations from textual descriptions, and such mapping is further leveraged by a facilitator model in ProteinDT to facilitate text-guided protein generation and editing. We will discuss the facilitator component in detail next.
ProteinFacilitator Model
The ProteinFacilitator model in ProteinDT aims at mapping a piece of text (a text prompt describing the properties of the target protein) to a protein representation that captures both the semantic information of the text and essential protein sequence patterns. This protein representation is then subsequently used for protein generation and editing. Such an explicit representation mapping from text modality to protein modality is inspired by the foundation models developed in vision [89] and chemistry [97], which has shown to improve the generalization quality (e.g., image diversity).
Concretely, the ProteinFacilitator model explicitly produces the protein sequence representation from the text representation, and we use a Gaussian distribution to model such a process. We introduce an alignment module $f_{\text{ProteinFacilitator}}(\cdot)$ and then we estimate the ProteinFacilitator distribution using a conditional Gaussian as $p({\bm{z}}_{\text{p}}|{\bm{z}}_{\text{t}})=\mathcal{N}({\bm{z}}_{\text{p}};f_{\text{ProteinFacilitator}}({\bm{z}}_{\text{t}}),I)$. To minimize its likelihood, the objective is essentially the representation reconstruction (regression) task [72]:
$\mathcal{L}_{\text{ProteinFacilitator}}=\|f_{\text{ProteinFacilitator}}({\bm{z}}_{\text{t}})-{\bm{z}}_{\text{p}})\|^{2},$ (6)
where $f_{\text{ProteinFacilitator}}(\cdot)$ is parameterized by a multi-layer perception. The produced protein sequence representation, i.e., $\tilde{\bm{z}}_{\text{p}}=f_{\text{ProteinFacilitator}}({\bm{z}}_{\text{t}})$, will be treated as the input condition, i.e., ${\bm{z}}_{\text{c}}=\tilde{\bm{z}}_{\text{p}}$, for the decoder model, as will be discussed in the next section. It is worth noting that, we also directly use the text sequence representation encoded by ProteinCLAP as the condition, namely ${\bm{z}}_{\text{c}}={\bm{z}}_{\text{t}}$. In the Results Section, we empirically verify the effectiveness of introducing the ProteinFacilitator model on the text-to-protein design task.
Protein Decoding
With the condition ${\bm{z}}_{\text{c}}$ generated by the aforementioned ProteinFacilitator, a decoder model is applied to conduct conditional generative modeling for protein generation, i.e., $p({\bm{x}}_{\text{p}}|{\bm{z}}_{\text{c}})$. In detail, during the training stage of ProteinDT, we directly take ${\bm{z}}_{\text{c}}={\bm{z}}_{\text{p}}$ to train the decoder in a teacher-forcing manner. In the next two sections, we introduce two types of conditional generative models for protein sequence generation: autoregressive (AR) and diffusion generative (ProteinDiff).
An Autoregressive Decoder
The training objective of AR is the next-token prediction (in Equation 1). As for model architecture, we adopt the Transformer for AR modeling [27], which is one of the most expressive architectures for various AR tasks, such as text-to-text [76, 77], text-to-speech [78, 79], and text-to-image generation [93]. More concretely, here we train a T5-Base model from scratch [63]. For conditional protein generation, the T5 decoder will generate the protein sequence token-by-token, where each generated token is conditioned on both the input condition and the previously decoded tokens. An illustration is shown in Supplementary C.
A Discrete Diffusion Decoder
We additionally investigate the protein sequence decoder employing the diffusion model within the ProteinDT framework. The diffusion generative model has been prevalent to solve the density estimation of vision tasks[66, 64, 81, 67]. More recently, the multinomial diffusion model [82, 136] has been proposed for modeling and generating discrete data such as textual sequence. We adopt the multinomial diffusion for protein sequence generation and call it ProteinDiff. The key component of ProteinDiff consists of a forward process that systematically introduces noise to the protein sequence ${\bm{x}}_{\text{p}}$ and a reverse process that performs denoising to reconstruct the original protein sequence. An illustration is shown in Supplementary C.
Forward process. This process adds noise to the input at each of the $T$ timesteps to convert $\tilde{\bm{x}}_{0}\to\tilde{\bm{x}}_{T}$, where $\tilde{\bm{x}}_{0}={\bm{x}}_{\text{p}}$ is the protein sequence and $\tilde{\bm{x}}_{T}$ is a random noise following a predefined distribution as will be introduced next. Specifically, given a condition representation ${\bm{z}}_{\text{c}}$, the forward transition of ProteinDiff at timestep $t$ is defined as:
$q(\tilde{\bm{x}}_{t}|\tilde{\bm{x}}_{t-1},{\bm{z}}_{\text{c}})=\text{Categorical}(\tilde{\bm{x}}_{t};p=\tilde{\bm{x}}_{t-1}Q_{t}),$ (7)
where $\tilde{\bm{x}}_{t}$ is the diffused protein sequence and $Q_{t}$ is the transition matrix at timestep $t$, respectively. We further take the absorbing transition [136] when adding noises: $Q_{t}=(1-\beta_{t})I+\beta\mathbbm{1}e_{m}^{T}$, where $\beta_{t}$ is the variance scheduled at timestep $t$ and $e_{m}$ is a one-hot vector with one on the absorbing state $m$ and zeros elsewhere. Specifically, we take 30 tokens as the vocabulary (including 26 alphabetic characters and 4 special tokens such as “[PAD]”), and the “[MASK]” token is used as the absorbing state $m$. Other variations of the transition matrix $Q_{t}$ are reserved for future exploration [136].
Objective and parameterization. To parameterize the above diffusion process, we adopt the following training tricks. Recent works [136, 84, 135] have empirically shown that the $\tilde{\bm{x}}_{0}$-paramtererization benefits more for the discrete data in diffusion modeling with reduced stochasticity. We therefore adopt such a strategy. That is, at each timestep $t$, we train a conditional transition network $p_{\theta}(\tilde{\bm{x}}_{0}|\tilde{\bm{x}}_{t},{\bm{z}}_{\text{c}})$ to predict the ground-truth protein sequence tokens. By taking the expectation over all $T$ timesteps, the objective function of ProteinDiff is the following for the masked tokens:
$\mathcal{L}_{\text{ProteinDiff}}=\mathbb{E}_{\tilde{\bm{x}}_{0}}\mathbb{E}_{t}\mathbb{E}_{\tilde{\bm{x}}_{t}|\tilde{\bm{x}}_{0}}\big{[}-\log p_{\theta}\big{(}\tilde{\bm{x}}_{0}|\tilde{\bm{x}}_{t},{\bm{z}}_{\text{c}}\big{)}\big{]}.$ (8)
The aforementioned transition network takes as input the condition representation ${\bm{z}}_{\text{c}}$ and the noised sequence $\tilde{\bm{x}}_{t}$ at timestep $t$, and outputs the predicted masked amino acids at timestep $t+1$. We consider two transition network architectures, a recurrent neural network (RNN) transition network and a Transformer transition network (BERT-Base) [26], for modeling.
Reverse process and decoding. The reverse process $\tilde{\bm{x}}_{T}\to\tilde{\bm{x}}_{0}$ in the diffusion model corresponds to the protein sequence sampling for protein generation in ProteinDiff. In detail, to generate a protein sequence, we first generate a diffused sequence $\tilde{\bm{x}}_{T}$ with all tokens being the absorbing state $m$. Then we consider two sampling strategies. (1) The simplified sampling [135]: for timestep $t$, we sample the mask positions and predict the masked tokens following the distribution $p_{\theta}(\tilde{\bm{x}}_{0}|\tilde{\bm{x}}_{t+1},{\bm{z}}_{\text{c}})$ for $\tilde{\bm{x}}_{t-1}$. (2) The weighted sampling [136]: for timestep $t$, the $p_{\theta}(\tilde{\bm{x}}_{t}|\tilde{\bm{x}}_{t+1})$ is indeed $p_{\theta}(\tilde{\bm{x}}_{0}|\tilde{\bm{x}}_{t+1},{\bm{z}}_{\text{c}})$ weighted by a posterior distribution (more details are in Supplementary C). We repeat such sampling process for $T$ timesteps, from $\tilde{\bm{x}}_{T}$ to $\tilde{\bm{x}}_{0}$, which yields protein sequence generation.
Acknowledgements
This project was partly done during Shengchao Liu’s internship at Nvidia and Ph.D. program at Mila-UdeM, and was supported in part by the Natural Sciences and Engineering Research Council (NSERC) Discovery Grant, the Canada CIFAR AI Chair Program, collaboration grants between Microsoft Research and Mila, Samsung Electronics Co., Ltd., Amazon Faculty Research Award, Tencent AI Lab Rhino-Bird Gift Fund, two NRC Collaborative R&D Projects, IVADO Fundamental Research Project grant PRF-2019-3583139727, and NSF award CHE 2226451.
Author Contributions Statement
S.L., Y.L., A.G., Y.Z., Z.X., W.N., A.R., C.X., J.T., H.G., and A.A. conceived and designed the experiments. S.L. performed the experiments. S.L., Y.L., and A.G. analyzed the data. S.L., Y.L., Z.L., and J.L. contributed analysis tools. S.L., Y.L., Z.L., A.G., C.X., H.G., and A.A. wrote the paper. C.X., J.T., H.G., and A.A. contributed equally to advising this project.
Competing Interests Statement
The authors declare no competing interests.
Supplementary Information
Appendix A Related Work
A.1 Multi-modality Learning
Recently, there has been a resurgent research interest in multi-modal learning in the machine learning community, exemplified by CLIP [86]. CLIP bridges the gap between image and language modalities by aligning the representation of the two via contrastive learning. This operation opens a novel research track on text-to-image (controllable) generation with natural language [87, 88, 89, 90, 91, 92, 93].
Multi-modality learning on small molecules. One active research line is utilizing external information to help augment the molecule representation learning, e.g., using a biological knowledge graph [94] or biomedical textual information [95, 96, 97] to augment the molecule representation learning. SGNN-EBM [94] and Text2Mol [95] focus on the supervised setting, while KV-PLM [96] aims at pre-training with a masked auto-encoding paradigm. MoMu [98] follows the setting of KV-PLM and further tests the text-to-molecule generation as in Text2Mol. MoleculeSTM [97] applies the contrastive learning paradigm to align the two modalities and conducts the text-guided molecule editing task. It reveals the superior understandability of natural language for finding novel biochemical concepts.
Multi-modality learning on macromolecules. For macromolecules like proteins, OntoProtein [99] first combines the protein sequence and textual description into a biological knowledge graph (KG), and the pretraining is done by the relation prediction on this biological KG. Chroma [100] is a parallel work that also conducts text-guided protein programmable editing, the ProCap model. Meanwhile, there are fundamental differences between ProCap and ProteinDT, as discussed in Table 4.
A.2 Protein Sequence and Structure Design
Generative protein sequence models and their applications were recently reviewed by [101]. Many recent protein sequence design models generate the protein sequence in an autoregressive manner [102, 103, 104, 105], i.e., amino-acid by amino-acid. Other protein sequence generation models incorporate autoencoders [106, 107, 108, 109, 110], generative adversarial networks [111, 112], or iterative masking with multiple sequence alignments [113].
Alternatively, reinforcement learning can also be used to generate sequences [114, 115, 116] or optimize sequences [117, 118, 119]. The reinforcement learning optimization setting resembles the ProteinDT text-guided protein editing task in that they have the same goal of improving an initial sequence. However, the reinforcement learning approach usually requires labels from experimental data unless the target property can be simulated [117]. In contrast, ProteinDT is a zero-shot approach requiring no experimental labels, only a suitable text prompt.
Other generative models include structure in the design process [120]. One approach is to start from a 3D protein backbone structure and design a compatible amino acid sequence [121, 122, 123, 124], which is known as inverse folding. Alternatively, there are also diverse techniques for designing 3D structures [125, 126, 127, 100, 128, 129, 130].
In Table 4, we provide a specific comparison between ProCap (Chroma) and ProteinDT.
Table 4: Comparison between ProCap (Chroma) and ProteinDT.
ProCap (Chroma) ProteinDT Data Structure ProCap or Chroma is focusing on the protein backbone structure. ProteinDT is focusing on the protein sequence. Dataset 45K text-backbone pairs from PDB and UniProt. 441K text-sequence pairs from SwissProtCLAP. Modeling ProCap builds a protein-to-text model (protein captioning), $p({\bm{x}}_{\text{t}}|{\bm{x}}_{\text{p},3D})$, where ${\bm{x}}_{\text{t}}$ is the textual description and ${\bm{x}}_{\text{p},3D}$ is the backbone (3D) structure. ProteinDT is mainly focusing on the text-to-protein model, $p({\bm{x}}_{\text{p}}|{\bm{x}}_{\text{t}})$, where ${\bm{x}}_{\text{p}}$ is the protein sequence. Inference ProCap utilizes classifier guidance [131], which is the MCMC sampling in a Bayes manner. The inference in ProteinDT depends on the decoder model. In the current version, autoregressive sampling and MCMC sampling are considered. Framework ProCap is specifically designed for a denoising diffusion generative model. ProteinDT is model-agnostic and thus is more flexible. Other deep generative models (energy-based model, variational auto-encoder, flow-based model, etc) can also be adopted.
Appendix B Pretraining Dataset SwissProtCLAP
We list the key statistics of the SwissProtCLAP dataset (as of date 2022.09.08) in Table 5. There are many redundant gene names, and we emphasize this by explicitly listing the number of protein identifiers and the number of unique gene names. The organism count is based on the number of unique NCBI Taxonomy identifiers.
Table 5: Statistics on SwissProtCLAP: the # text-protein pairs, # protein identifiers, # unique gene names, and # unique organisms.
# text-protein pairs # proteins # genes # organisms 441,013 441,013 327,577 13,339
Then we randomly select six text-protein pairs from SwissProtCLAP in Table 6 for illustration, so as to give a better intuition about the dataset contents.
Table 6: Examples of text-protein pairs in SwissProtCLAP.
Protein Sequence Text Sequence MSLNAEQKAKVVLEHGSSAHDTGSTEVQVALL
TLRINDLQKHFLEHKKDHHSRRGLLRMVSQRR
KLLDYLKKRNISKYTDLIQSLGLRK
One of the primary rRNA binding proteins, it binds directly to 16S rRNA where it helps nucleate assembly of the platform of the 30S subunit by binding and bridging several RNA helices of the 16S rRNA. Forms an intersubunit bridge (bridge B4) with the 23S rRNA of the 50S subunit in the ribosome. Part of the 30S ribosomal subunit. Forms a bridge to the 50S subunit in the 70S ribosome, contacting the 23S rRNA. Belongs to the universal ribosomal protein uS15 family.
MRATVGLVEAIGIRELRQHASRYLARVEAGEE
LGVTNKGRLVARLIPVQAAERSREALIESGVL
IPARRPQNLLDVTAEPARGRKRTLSDVLNEMR
Belongs to the phD/YefM antitoxin family.
LKPDEELQGPGGVLSRGYFVFRPRN
Stimulates uterine smooth muscle contraction and causes selective vasoconstriction. Belongs to the NmU family.
MARSLKKGPFVDHHLAKKVESAAGSKKPIKTW
SRRSMILPEMVGITIAVHNGKNHIPVLVNENM
VGHKLGEFAVTRTFKGHGGDKKSSR
Protein S19 forms a complex with S13 that binds strongly to the 16S ribosomal RNA. Belongs to the universal ribosomal protein uS19 family.
MAKVIIEIKNTVSGIKGRNLRTSIAVDGSAEL
DGDEGTLAGMVALLVLNKSQKIINESAHEAIE
ILKNDGVITSGRVTEMAVEKTCH
Expressed in the early phase of the viral replicative cycle. Expression of early genes is repressed by viral Repc (latency) and favored by viral Ner protein.
MSRTIFCTFLNKEADGLDFQLYPGELGKRIFN
EISKEAWGQWMAKQTMLINEKKLNTMNPDDRK
LLEQEMVRFLFEGHDVHIDGYTPPEK
Could be a mediator in iron transactions between iron acquisition and iron-requiring processes, such as synthesis and/or repair of Fe-S clusters in biosynthetic enzymes. Monomer. Belongs to the Fe(2+)-trafficking protein family.
Appendix C Architecture Details and Ablation Studies
We list the implementation details of ProteinDT, following the pretraining steps and downstream tasks.
C.1 ProteinCLAP Model
We use two BERT models to model the protein sequence and text sequence, respectively. Then we take the SciBERT [132] and the ProtBERT [102] as the starting checkpoint, followed by ProteinCLAP pretraining. We also add a prediction head to each of the BERT models since this approach has been found [133, 134] empirically useful in the contrastive learning setting. The key hyperparameters are listed below.
Table 7: Hyperparameter specifications for ProteinCLAP.
Hyperparameter Value epochs {5, 10} learning rate for text sequence representation {1e-5, 1e-6} learning rate for protein sequence representation {1e-5, 1e-6} latent dimension of prediction head {256} objective function {EBM-NCE, InfoNCE} temperature {0, 0.1 } batch size {8}
C.2 ProteinFacilitator Model
We consider using the Gaussian distribution as the ProteinFacilitator distribution in ProteinDT. The key hyperparameters are listed below.
Table 8: Hyperparameter specifications for ProteinFacilitator distribution.
Hyperparameter Value epochs {32} learning rate {1e-6} objective function {MSE} batch size {16}
C.3 Decoder Model
(a) Autoregressive model.
(b) Diffusion model.
Figure 4: The inference illustration for two decoder models (step 3 in ProteinDT). Figure 4(a) is an autoregressive (AR) model based on the Transformer. It generates the protein sequence token-by-token. Figure 4(b) is ProteinDiff, and a Transformer-based transition network is displayed. It first randomly samples a noised sequence, then conducts a denoising process using the transition network.
For the decoder, we consider two categories of models: the autoregressive (AR) and the discrete diffusion model (ProteinDiff). We list the pipeline in Figure 4, and the key hyperparameters are listed below.
Table 9: Hyperparameter specifications for decoder distribution.
model Hyperparameter Value Autoregressive epochs {10} learning rate {1e-4, 1e-5} ProteinDiff transition network architecture {RNN, Transformer} epochs {10} learning rate {1e-4, 1e-5} hidden dimension {16, 32}
Inference for AR
For the autoregressive (AR) model, we consider two variants in inference: the greedy search and the beam search.
Inference for ProteinDiff
The reverse process $\tilde{\bm{x}}_{T}\to\tilde{\bm{x}}_{0}$ in the diffusion model corresponds to the protein sequence sampling for protein generation in ProteinDiff. In detail, to generate a protein sequence, we first generate a diffused sequence ${\bm{x}}_{T}$ with all tokens being the “[MASK]” token. Then we consider two sampling strategies, as below.
(1) The simplified sampling [135]
For timestep $t$, we sample the mask positions with probability $1/t$ and predict the masked tokens following the distribution $p_{\theta}(\tilde{\bm{x}}_{0}|\tilde{\bm{x}}_{t+1},{\bm{z}}_{\text{c}})$. This is an approximation sampling, yet works quite well on discrete data, like quantized codes.
(2) The weighted sampling [136]
For timestep $t$, the $p_{\theta}(\tilde{\bm{x}}_{t}|\tilde{\bm{x}}_{t+1},{\bm{z}}_{\text{c}})$ is indeed $p_{\theta}(\tilde{\bm{x}}_{0}|\tilde{\bm{x}}_{t+1},{\bm{z}}_{\text{c}})$ weighted by a posterior distribution. More concretely, first, with the predefined transition matrix and forward process function, we can get the posterior $q(\tilde{\bm{x}}_{t}|\tilde{\bm{x}}_{t+1},\tilde{\bm{x}}_{0})$ in a closed-form as:
$q(\tilde{\bm{x}}_{t}|\tilde{\bm{x}}_{t+1},\tilde{\bm{x}}_{0})=\frac{q(\tilde{\bm{x}}_{t+1}|\tilde{\bm{x}}_{t})\cdot q(\tilde{\bm{x}}_{t}|\tilde{\bm{x}}_{0})}{q(\tilde{\bm{x}}_{t+1}|\tilde{\bm{x}}_{0})}=\text{Categorical}\Big{(}\tilde{\bm{x}}_{t};p=\frac{\tilde{\bm{x}}_{t+1}Q_{t+1}^{T}\odot\tilde{\bm{x}}_{0}\bar{Q}_{t}}{\tilde{\bm{x}}_{0}\bar{Q}_{t+1}\tilde{\bm{x}}_{t+1}^{T}}\Big{)},$ (9)
where $\bar{Q}_{t}=\prod_{s=1}^{t}Q_{s}$, and $\odot$ is the element-wise product. Then we can get the logits $\tilde{\bm{x}}_{0}$ using the transition network (same as the simplified sampling), i.e., $p_{\theta}(\tilde{\bm{x}}_{0}|\tilde{\bm{x}}_{t},{\bm{z}}_{\text{c}})$. Thus the weighted sampling for $\tilde{\bm{x}}_{t}$ is:
$p_{\theta}(\tilde{\bm{x}}_{t}|\tilde{\bm{x}}_{t+1},{\bm{z}}_{\text{c}})=\sum_{\bar{\bm{x}}_{0}}q(\tilde{\bm{x}}_{t}|\tilde{\bm{x}}_{t+1},\bar{\bm{x}}_{0})\cdot p_{\theta}(\bar{\bm{x}}_{0}|\tilde{\bm{x}}_{t+1},{\bm{z}}_{\text{c}}),$ (10)
where $\bar{\bm{x}}_{0}$ is the one-hot representation of $\tilde{\bm{x}}_{0}$.
The effect of sampling strategies.
Our ablation studies, detailed in Table 10, indicate that there isn’t a significant disparity between the two sampling methods. Notably, employing RNN with a simplified sampling method and utilizing the backbone yields preferable outcomes, whereas Transformer-based ProteinDiff performs better with weighted sampling. Hence, for the sake of simplicity, we present the results exclusively using simplified sampling. This finding holds promise for future explorations in this domain.
Table 10: Ablation studies on using simplified and weighted sampling in ProteinDiff for text-to-protein-generation. The retrieval accuracies (%) are reported with different $T$-choose-1 tasks.
Model $T=4$ $T=10$ $T=20$ ProteinDiff-RNN simplified sampling 40.50 21.50 15.00 weighted sampling 33.50 17.00 13.00 ProteinDiff-Transformer simplified sampling 51.50 25.00 13.50 weighted sampling 53.50 28.50 21.00
Appendix D Task Specification
D.1 Downstream Task: Text-to-Protein Generation
As discussed in the main manuscript, one of the most important tricks is whether to not to use ProteinFacilitator. In this section, we would like to discuss other important tricks and hyperparameters for text-to-protein generation.
D.1.1 Sequence Length
For the AR model, there is an end-of-sequence (“EOS”) token explicitly telling the end of generated sequences. For ProteinDiff, we either take the first padding token (“PAD”) as the end of the protein sequence, or we ignore all the padding tokens in the generated protein sequences.
D.1.2 Selection of Generated Protein Sequences
Note that we also incorporate ProteinCLAP to help augment the quality of generated protein sequences. Concretely, for each textual description, we generate 16 protein sequences, then we select the one with the highest similarity with the textual description. We conduct this for all the baselines and ProteinDT.
D.2 Downstream Task: Text-guided Protein Editing
Zero-shot text-guided protein editing is one of the most important tasks we design. Due to the space limitation in the main manuscript, we would like to provide more details in this section.
D.2.1 Editing Models
We cover two editing methods: latent interpolation and latent optimization. Latent interpolation does not require extra training, yet latent optimization requires us to train a decoder $g$. We list the detailed comparison between the two in Table 11. Specifically in latent optimization, we consider a recurrent neural network (RNN) model, and the hyperparameters are listed in Table 12.
Determining the appropriate decoder in latent optimization remains an unresolved query. Within ProteinDT, our initial step in this avenue of research involves the utilization of RNN, recognized as a highly efficient and proficient sequential model. Empirical evidence substantiates the commendable performance of RNN as a decoder in executing protein design tasks. It is imperative to note that the investigation into more sophisticated decoder architectures is deferred for future exploration.
Table 11: Comparison of key modules between latent interpolation and latent optimization.
Latent Interpolation Latent Optimization Representation Sequence-level Token-level Decoding Token-by-token Token Prediction Learning No Learning Learning Required Text-guidance SLERP Interpolation Similarity
Table 12: Hyperparameter specifications for latent optimization decoder (RNN).
Hyperparameter Value epochs {10} learning rate {1e-3, 5e-3} RNN layer {2, 3}
D.2.2 Dataset of Editing Tasks
In this section, we would like to provide more details on the dataset of each editing task. All the data preprocessing scripts can be found in the GitHub repository.
Structure editing. We follow prior work on protein secondary structure prediction [137]. The oracle evaluator (as will be discussed next) is pretrained on the training set, so we take the test set as input protein sequences (data splitting from the original paper). Three distinct test sets are available, and in accordance with [138], we opt for the “cb513” subset, comprising 513 protein sequences. This holds for the editing tasks on both alpha-helix and beta-sheet.
Stability editing. Rocklin et al. assessed protein stability with a high-throughput assay based on protease resistance [137]. Cells were treated with multiple concentrations of a protease. A protein’s measured ability to tolerate the protease was normalized by its estimated tolerance in an unfolded state to generate a stabilty score. Therefore, larger positive stability scores represent more stable proteins. Negative scores mean that the unfolded state was estimated to be more resistant to the protease treatment than what was measured experimentally for that protein.
Although Rocklin et al. and the TAPE [139] stability dataset derived from this study contain stability measurements for multiple rounds of iterative protein design as well as saturation mutagenesis experiments for designed and natural proteins, we focus on two natural proteins, villin HP35 and Pin1 WW domain. In this controlled setting, we increase the likelihood that the regressor will be able to estimate reasonable stability scores for the output edited sequences. If we were to instead train the regressor on all available sequences and edit a wide variety of input sequences, we may be less confident in the regressor’s ability to score those generated sequences accurately.
We extract the villin saturation mutagenesis stability data from the TAPE dataset, which places all of these sequences in the test split, by selecting all 631 instances with the topology set to villin. The saturation mutagenesis data contains the wild type villin HP35 sequence and all 18 possible single amino acid mutations at each position in the sequence, excluding cysteine (C). We remove the five-character prefix GSSGS and three-character suffix GSS from the protein sequences before training the regressor and providing the sequeces as input to the editing model. The prefix and suffix represent padding residues added during library design [137] and are not part of the 35 amino acids long villin sequence. Similarly, we extract Pin1 stability data from the TAPE dataset. We select all 703 instances with the topology set to Pin1. We remove the one-character prefix G and three-character suffix GSS, leading to 39 amino acids in the Pin1 sequences. We do not remove prefix or suffix sequences for the TAPE stability prediction task above.
Peptide binding editing. We follow recent work of estimating the protein-ligand binding affinity and dynamics [140]. Here we extract the protein complexes with a peptide as the ligand. On the other hand, we also need the textual description of the protein sequences. To achieve this, we keep the protein-ligand complexes wherein the proteins exist with the SwissProtCLAP. Following the filtration process, this results in a total of 124 peptide sequences.
D.2.3 Sequence Length
For the length of the generated sequences, we maintain consistency with the strategy used for text-to-protein generation. However, an additional constraint is introduced: the maximum sequence length must fall within a specific range relative to the input protein sequences. To do so, we dynamically establish the maximum sequence length for generated proteins by adopting the maximum protein sequence length from each batch of protein sequences. Namely, we take the maximum protein sequence length for each input batch as the upper limit for the generated protein sequences.
Furthermore, in the context of the stability editing and peptide binding editing tasks, the input protein sequences are relatively compact, reaching a maximum of 39 and 15 amino acids, respectively. However, for the structure editing task, the computational burden incurred during the folding evaluation is considerable. Consequently, we take all the protein sequences with length over 512 as invalid proteins.
D.2.4 Selection of Generated Protein Sequences
Additionally, we incorporate ProteinCLAP to enhance the quality of generated protein sequences. In alignment with text-to-protein generation, we generate 16 protein sequences for each input protein sequence and text prompt. Subsequently, we select the sequence exhibiting the highest similarity with either the input description or input protein sequence. This selection methodology is applied within ProteinDT. However, it differs from the baselines (ChatGPT and Galactica), as they amalgamate the input protein sequence and text prompt into a singular sentence, and such a design framework does not support the post-sampling selection process outlined above. It’s crucial to note that the primary distinction between editing and text-to-protein generation lies in the uniformity of inputs for the latter across all baselines and ProteinDT.
D.2.5 Evaluation Function
Next, we would like to list more details on the oracle evaluation function of different editing tasks.
Structure editing. One oracle evaluation function we considered is a protein sequence evaluator for secondary structure prediction (as discussed in the property prediction downstream below). More concretely, we are using the ProteinDT-ProteinCLAP-EBM-NCE, which is the best predictor of SS-Q3 in our evaluations.
Stability editing. The oracle evaluation function we considered is the optimal protein sequence evaluator for stability (as discussed in the property prediction downstream below). More concretely, we are using ProtBERT, which is the best predictor of Stability in our evaluations.
Peptide binding editing. For the peptide editing, we first train an oracle model on the protein-peptide binding datasets from PDB. The geometric model satisfies the SE(3)-equivariance using the vector frame basis [141], which is one of the state-of-the-art models on protein structure modeling. Then we adopt GNINA as one of the oracle evaluation metrics [142], which is a docking software that uses an ensemble of convolutional neural networks as the scoring function. For each conformation, it simultaneously predicts the pose score (probability that the current pose resembles the actual pose) and binding affinity. An ideal docking conformation should have a high affinity under the pose score of 1. Thus, the final docking score is calculated as the product of the pose score and binding affinity, as the evaluation of the current docking conformation quality.
Satisfactory function. The oracle evaluation functions listed above are for evaluating one protein sequence (either input protein sequence or output protein sequence). The next step is to calculate the satisfactory hit ratio (%), which measures the ratio of the output score being higher or lower than the input score, as indicated in the text prompt.
D.2.6 Text Prompt in Text-guided Protein Editing
In this section, we add details of the text prompts for 10 editing tasks with ProteinDT and two baselines.
Table 13: Prompts for structure editing. The curly brackets “{}” represent the input protein sequence.
Task Method Prompt more alpha helices Galactica Given an input amino acid sequence [START_AMINO]{}[END_AMINO]. Can you modify the amino acid sequence to have more alpha helices in the secondary, which needs to be similar to the input sequence? [START_AMINO] ChatGPT Given an input amino acid sequence {}. Can you modify the amino acid sequence to have more alpha helices in the secondary structure, which needs to be similar to the input sequence (just AA sequence, and no explanation)? ProteinDT modify the amino acid sequence to have more alpha helices in the secondary structure less alpha helices Galactica Given an input amino acid sequence [START_AMINO]{}[END_AMINO]. Can you modify the amino acid sequence to have less alpha helices in the secondary, which needs to be similar to the input sequence? [START_AMINO] ChatGPT Given an input amino acid sequence {}. Can you modify the amino acid sequence to have less alpha helices in the secondary structure, which needs to be similar to the input sequence (just AA sequence, and no explanation)? ProteinDT modify the amino acid sequence to have less alpha helices in the secondary structure more beta sheets Galactica Given an input amino acid sequence [START_AMINO]{}[END_AMINO]. Can you modify the amino acid sequence to have more beta sheets in the secondary, which needs to be similar to the input sequence? [START_AMINO] ChatGPT Given an input amino acid sequence {}. Can you modify the amino acid sequence to have more beta sheets in the secondary structure, which needs to be similar to the input sequence (just AA sequence, and no explanation)? ProteinDT modify the amino acid sequence to have more beta sheets in the secondary structure less beta sheets Galactica Given an input amino acid sequence [START_AMINO]{}[END_AMINO]. Can you modify the amino acid sequence to have less beta sheets in the secondary, which needs to be similar to the input sequence? [START_AMINO] ChatGPT Given an input amino acid sequence {}. Can you modify the amino acid sequence to have less beta sheets in the secondary structure, which needs to be similar to the input sequence (just AA sequence, and no explanation)? ProteinDT modify the amino acid sequence to have less beta sheets in the secondary structure
Table 14: Prompts for stability editing. The following are used for both villin and Pin1, and the curly brackets “{}” represent the input protein sequence.
Task Method Prompt higher stability Galactica Given an input amino acid sequence [START_AMINO]{}[END_AMINO]. Can you modify the amino acid sequence to have higher stability, which needs to be similar to the input sequence? [START_AMINO] ChatGPT Given an input amino acid sequence {}. Can you modify the amino acid sequence to have higher stability, which needs to be similar to the input sequence (just AA sequence, and no explanation)? ProteinDT modify the amino acid sequence to have higher stability lower stability Galactica Given an input amino acid sequence [START_AMINO]{}[END_AMINO]. Can you modify the amino acid sequence to have lower stability, which needs to be similar to the input sequence? [START_AMINO] ChatGPT Given an input amino acid sequence {}. Can you modify the amino acid sequence to have lower stability, which needs to be similar to the input sequence (just AA sequence, and no explanation)? ProteinDT modify the amino acid sequence to have lower stability
Table 15: Prompts for peptide binding editing. The curly brackets “{}” represent each binding task’s input peptide sequence and protein property description accordingly. The detailed descriptions of each protein target can be found in the GitHub repository.
Task Method Prompt higher affinity Galactica Given an input peptide amino acid sequence [START_AMINO]{}[END_AMINO] and a target protein. The target protein satisfies the following property. {} Can you modify the peptide amino acid sequence to have higher binding affinity with the target protein, which needs to be similar to the input sequence? [START_AMINO] ChatGPT Given an input peptide amino acid sequence {} and a target protein. The target protein satisfies the following property. {} Can you modify the peptide amino acid sequence to have higher binding affinity with the target protein, which needs to be similar to the input sequence (just AA sequence, and no explanation)? ProteinDT modify the peptide amino acid sequence to have higher binding affinity with the target protein. The target protein satisfies the following property. {} lower affinity Galactica Given an input peptide amino acid sequence [START_AMINO]{}[END_AMINO] and a target protein. The target protein satisfies the following property. {} Can you modify the peptide amino acid sequence to have lower binding affinity with the target protein, which needs to be similar to the input sequence? [START_AMINO] ChatGPT Given an input peptide amino acid sequence {} and a target protein. The target protein satisfies the following property. {} Can you modify the peptide amino acid sequence to have lower binding affinity with the target protein, which needs to be similar to the input sequence (just AA sequence, and no explanation)? ProteinDT modify the peptide amino acid sequence to have lower binding affinity with the target protein. The target protein satisfies the following property. {}
D.3 Downstream Task: Property Prediction
The detailed specifications for the six TAPE [139] tasks can be found in Table 16.
Table 16: Dataset specification.
dataset type # train # val #test # labels
SS-Q3 token-level classification 8,678 2,170 649 3
SS-Q8 token-level classification 8,678 2,170 649 8
Contact token-level (token pair) classification 24,651 224 40 2
Homology sequence-level classification 12,312 736 1,976 1,195
Fluorescence sequence-level regression 21,446 5,362 27,217 -
Stability sequence-level regression 53,614 2,512 12,851 -
We follow the same hyperparameters in OntoProtein [99], as shown in Table 17.
Table 17: Hyperparameter specifications for protein property prediction.
Task epoch batch size warmup ratio learning rate frozen BERT Optimizer SS-Q3 5 32 0.08 3e-5 False AdamW SS-Q8 5 32 0.08 3e-5 False AdamW Contact 10 8 0.08 3e-5 False AdamW Homology 10 64 0.08 3e-5 False AdamW Fluorescence 25 64 0 3e-5 True AdamW Stability 5 32 0.08 3e-5 False AdamW

### v4

@@
 To train ProteinDT, we construct a large dataset, SwissProtCLAP, with 441K text and protein pairs.
-We quantitatively verify the effectiveness of ProteinDT on three challenging tasks: (1) over 90% accuracy for text-guided protein generation; (2) best hit ratio on 10 zero-shot text-guided protein editing tasks; (3) superior performance on four out of six protein property prediction benchmarks.
+We quantitatively verify the effectiveness of ProteinDT on three challenging tasks: (1) over 90% accuracy for text-guided protein generation; (2) best hit ratio on 12 zero-shot text-guided protein editing tasks; (3) superior performance on four out of six protein property prediction benchmarks.
 Machine learning (ML) has recently shown profound potential for protein discovery.
 These ML tools have been quickly adapted as auxiliary and accelerating roles in scientific pipelines, including but not limited to protein engineering [1], structure prediction [2], structure reconstruction [3], and inverse folding [4].
-To this end, most existing such methods merely model the intrinsic properties of protein sequences and structures [101, 6, 102, 8, 9, 10, 11].
+To this end, most existing such methods merely model the intrinsic properties of protein sequences and structures [5, 6, 7, 8, 9, 10].
 Motivated by recent breakthroughs of foundation models [11, 12, 13, 14], approaches in the computational chemistry domain that utilize textual descriptions of drugs in addition to their intrinsic chemical and structural information have proved to be effective in small-molecule drug discovery [15, 16, 17, 18, 19].
@@
 Both editing methods inject the text modality and the learned representation is used for protein generation accordingly.
-We observe that ProteinDT can reach consistently the best hit ratio on 10 editing tasks, ranging from structure- and stability-aware tasks, to peptide-binding editing tasks.
+We observe that ProteinDT can reach the overall best hit ratio on 12 editing tasks, ranging from structure- and stability-aware tasks, to peptide-binding editing tasks.
 Additional qualitative results offer further compelling evidence supporting the validity of ProteinDT.
 The last task is protein property prediction, aiming to evaluate the robustness and generalization ability of the learned representation from ProteinCLAP.
-Compared to six state-of-the-art protein sequence representation methods, ProteinDT can obtain consistently superior performance on four of six benchmark tasks.
-Figure 1: Pipeline for ProteinDT framework (a-c) and downstream tasks (d-f).
+Compared to six state-of-the-art protein sequence representation methods, ProteinDT can obtain superior performance on four of six benchmark tasks.
+Figure 1: Pipeline for ProteinDT pretraining framework (a-c) and downstream tasks (d-f).
 (a) ProteinCLAP, a contrastive learning paradigm, aligns the representation space of the text and protein sequence modalities.
@@
 (d) Downstream text-to-protein generation task.
-(e) Downstream text-guided protein design task.
+(e) Downstream text-guided protein editing task.
 (f) Downstream protein property prediction task.
@@
 There are in total 20 different amino acids in proteins, represented with 20 characters, as shown in Figure 1.
-For the protein sequence representation ${\bm{z}}_{\text{p}}=f_{\text{p}}({\bm{x}}_{\text{p}})$, we use BERT as the encoder [26].
+For the protein sequence-level representation ${\bm{z}}_{\text{p}}=f_{\text{p}}({\bm{x}}_{\text{p}})$, we use BERT as the encoder [25].
 It is one of the most advanced Transformer models for modeling sequential data [26].
@@
 Text sequence ${\bm{x}}_{\text{t}}$ and representation ${\bm{z}}_{\text{t}}$.
-The textual descriptions ${\bm{x}}_{\text{t}}$ are formulated in the form of sequences of tokens, and we also adopt the BERT model as the encoder ${\bm{z}}_{\text{t}}=f_{\text{t}}({\bm{x}}_{\text{t}})$.
+The textual descriptions ${\bm{x}}_{\text{t}}$ are formulated in the form of sequences of tokens, and we also adopt the BERT model as the encoder to get the sequence-level representation ${\bm{z}}_{\text{t}}=f_{\text{t}}({\bm{x}}_{\text{t}})$.
 We further adapt the pretrained SciBERT [29], which was pretrained on the full-text computer science and biomedical papers from Semantic Scholar [30].
@@
 Last but not least, ProteinDT subsequently utilizes a conditional generative model to decode the formed protein representation into the protein sequence.
-We propose one autoregressive model (AR) with Transformer and two diffusion models (ProteinDiff).
-More details of three pretraining steps can be found in the Methods section.
-With such a ProteinDT framework, we can explore a broad range of challenging protein design tasks, as will be illustrated below.
+We propose one autoregressive model (AR) and two diffusion models (ProteinDiff).
+More details of the three pretraining steps can be found in the Methods section.
 Downstream Task: Text-to-Protein Generation
@@
 (b-c) Two methods for text-guided protein editing: latent interpolation and latent optimization.
-(d-f) Visualization for evaluation of text-guided protein editing.
-For three types of editing tasks, different evaluation metrics (marked in red) are applied accordingly.
+(d-e) Visualization for evaluation of text-guided protein editing.
+For four types of editing tasks, different evaluation metrics (marked in red) are applied accordingly.
 Table 1: Retrieval accuracy (%) for text-to-protein generation. It is a multiple-choice task ($T$ options), measuring the consistency between the input text prompt and the generated protein sequence.
-Galactica ChatGPT ProteinDT (Ours) w/o ProteinFacilitator w/ ProteinFacilitator AR AR AR ProteinDiff AR ProteinDiff Transformer RNN Transformer Transformer RNN Transformer T = 4 51.5 38.5 49.00 24.00 35.50 97.00 40.50 51.50 T = 10 29.0 23.0 27.00 10.50 17.50 91.00 21.50 25.00 T = 20 19.0 15.5 20.00 5.50 9.50 83.50 15.00 13.50
+Galactica ChatGPT ProteinDT (Ours)
+w/o ProteinFacilitator w/ ProteinFacilitator
+AR AR AR ProteinDiff AR ProteinDiff
+Transformer RNN Transformer Transformer RNN Transformer
+T = 4 51.5 38.5 49.00 24.00 35.50 97.00 40.50 51.50
+T = 10 29.0 23.0 27.00 10.50 17.50 91.00 21.50 25.00
+T = 20 19.0 15.5 20.00 5.50 9.50 83.50 15.00 13.50
 Experiment setting.
@@
 For evaluation, we adopt the well-trained ProteinCLAP to obtain the retrieval accuracy: each protein sequence is compared to the input text prompt and $T-1$ randomly chosen texts using ProteinCLAP, and a match is successful if the actual text-sequence pair ranks the highest among all comparisons.
-The evaluation pipeline is shown in Figure 2.
+The evaluation pipeline is shown in Figure 2, and please check Supplementary D for more details.
 Baselines.
@@
 Observation 2.
-For the three decoder variants of ProteinDT, AR-Transformer can beat two diffusion models (ProteinDiff).
+For the three decoder variants of ProteinDT, AR-Transformer can beat the two diffusion models (ProteinDiff).
 Similar findings are also reported in recent studies on text-to-text generation [34, 35].
@@
 Latent interpolation.
-We have two representations for the text prompt and input protein as ${\bm{z}}_{\text{t}}$ and ${\bm{z}}_{\text{p}}$, respectively.
+Latent interpolation aims at getting the latent codes by directly interpolating two endpoint representations [12, 13, 36], where the latent codes will be used for sampling data points through the decoder.
+In this setting, we have two representations for the text prompt and input protein as ${\bm{z}}_{\text{t}}$ and ${\bm{z}}_{\text{p}}$, respectively.
 Then we apply the spherical linear interpolation (SLERP) [13], which yields an intermediate representation $\tilde{\bm{z}}_{\text{c}}=\text{SLERP}({\bm{z}}_{\text{t}},{\bm{z}}_{\text{p}},\theta)$, and the output protein is generated from decoder, as $\tilde{\bm{x}}_{\text{p}}\sim p({\bm{x}}_{\text{p}}|\tilde{\bm{z}}_{\text{c}})$.
@@
 Latent optimization.
-We also propose using latent optimization [90, 97] for text-guided protein editing.
-The high-level idea is to learn a token-level latent code ${\bm{z}}_{w}$ that is close to both the text and protein sequence-level representation, as ${\bm{z}}_{w}=\arg\min(\lambda\|P({\bm{z}}_{w})-{\bm{z}}_{\text{t}}\|^{2}+(1-\lambda)\|P({\bm{z}}_{w})-{\bm{z}}_{\text{p}}\|^{2})$, where $P$ is the pooling function, and $\lambda$ controls the closeness to the target prompt.
+We also propose using latent optimization [14, 37, 18] for text-guided protein editing.
+The high-level idea is to directly optimize a latent code in the representation space and then decode it to get the sampled data.
+Specifically in ProteinDT, it learns a token-level latent code ${\bm{z}}_{w}$ that is close to both the text and protein sequence-level representation, as ${\bm{z}}_{w}=\arg\min(\lambda\|P({\bm{z}}_{w})-{\bm{z}}_{\text{t}}\|^{2}+(1-\lambda)\|P({\bm{z}}_{w})-{\bm{z}}_{\text{p}}\|^{2})$, where $P$ is the pooling function, and $\lambda$ controls the closeness to the target prompt.
 Then, the learned ${\bm{x}}_{w}$ are mapped to the protein sequence using a pretrained decoder $g$, i.e., $\tilde{\bm{x}}_{\text{p}}=g({\bm{z}}_{w})$.
@@
 Editing tasks and evaluation metrics.
-To fully verify the effectiveness of ProteinDT, we design 10 zero-shot text-guided protein editing tasks.
-We further categorize them into three settings as detailed below.
+To fully verify the effectiveness of ProteinDT, we design 12 zero-shot text-guided protein editing tasks.
+We further categorize them into four settings as detailed below.
 For editing evaluation, we employ the satisfactory hit ratio (%), which measures the ratio of the output score being higher or lower than the input score, as indicated in the text prompt.
-The scores are provided by oracle evaluators on each protein sequence, as discussed below.
-To mitigate bias resulting from different oracle evaluators, we adopt two distinct types of evaluators for each task.
+The scores are provided by evaluators on each protein sequence, as discussed below.
+To mitigate bias resulting from different evaluators, we adopt two distinct types of evaluators for each task (except region editing).
 Certain evaluations require proteins’ 3D structures, for which we employ the ESMFold algorithm [7].
@@
 The prompts are in the format of “modify the amino acid sequence to have more (less) alpha helices” and “modify the amino acid sequence to have more (less) beta sheets”.
-We adopt two methods to measure the secondary structures for each amino acid: (1) A strong protein sequence evaluator for secondary structure prediction (as will be discussed in the property prediction downstream below); (2) Define Secondary Structure of Proteins (DSSP) algorithm for assigning secondary structure to the amino acids [37].
+We adopt two methods to measure the secondary structures for each amino acid: (1) A protein sequence evaluator for secondary structure prediction (as will be discussed in the property prediction downstream below); (2) Define Secondary Structure of Proteins (DSSP) algorithm for assigning secondary structure to the amino acids [38].
+(2) Region editing.
+The second editing task is region editing.
+The predicted Local Distance Difference Test (pLDDT) is a per-residue confidence metric used to estimate the reliability of the predicted 3D structure of a protein sequence.
+We consider proteins from this paper [39], which reveals the relation between ordered/disordered regions and pLDDT.
+Ordered regions most frequently show higher pLDDT, while the disordered regions are reflected with low pLDDT.
+The prompts are in the format of “modify the amino acid sequence to have more ordered/disordered regions”.
+We adopt the pLDDT from ESMFold as the evaluation metric [7].
 (3) Stability editing.
-Then we consider four types of stability editing tasks.
+Then we consider four types of thermodynamic stability editing tasks.
 Two proteins are considered [40]: villin HP35 and Pin1 WW domain.
 The prompts are in the format of “modify the amino acid sequence to have higher (lower) stability”.
-We consider two evaluators: (1) A strong protein sequence evaluator for stability (as will be discussed in the property prediction downstream below); (2) The predicted local distance difference test (pLDDT) score measuring the confidence for folding, which has been used as a proxy for stability [2].
+We consider two evaluators: (1) A protein sequence evaluator for stability (as will be discussed in the property prediction downstream below); (2) Rosetta to quantify the energy after structure prediction, in Rosetta Energy Units (REU) [41, 42].
+The score function is REF2015 [43].
 (4) Peptide binding editing.
@@
 Here we take 154 pairs of peptide-protein binding pairs from Protein Data Bank (PDB) [44], where the proteins have textual descriptions from UniProt in SwissProtCLAP.
-Then, the task is to make the input peptide with higher/lower binding affinity with respect to the target protein (textual description).
-For scoring with the oracle evaluator, we consider two options: (1) A pretrained deep learning model for binding affinity prediction [140]; (2) GNINA to calculate the docking score before and after editing [142].
+Then, the task is to make the input peptide have higher/lower binding affinity with respect to the target protein (textual description).
+For scoring with the evaluator, we consider two options: (1) A pretrained deep learning model for binding affinity prediction [45]; (2) GNINA to calculate the docking score before and after editing [46].
 Baselines.
@@
 The main results are in Table 2.
-We can observe that two variants of ProteinDT (latent interpolation and latent optimization) generally perform better on all 10 tasks and 20 evaluation methods.
+We can observe that two variants of ProteinDT (latent interpolation and latent optimization) generally perform better on 12 tasks and 21 evaluation methods.
 However, for the stability editing tasks, latent interpolation methods tend to favor lower stability, while latent optimization can reach a good balance between higher and lower stability tasks.
-This suggests that latent optimization is in general a preferable method for protein editing.
+This suggests that latent optimization is, in general, a preferable method for protein editing.
 Despite being a generic foundation model, ChatGPT is occasionally the second-best performer on some of the stability editing tasks.
@@
 Figure 3 (a) and (d) show the structures of two input protein sequences.
-Then ProteinDT adopts the prompts with key words “more $\alpha$-sheets” and “more $\beta$-strands”, the edited protein structures (edited sequences folded with ESMFold) are shown in Figure 3 (b) and (e).
-Similarly for Figure 3 (c) and (f) when using prompts like “less $\alpha$-sheets” and “less $\beta$-strands”.
-For peptide editing illustration, we take PDB 3IQI, Haemophilus influenzae O-acetylserine sulfhydrylase in complex with the peptide MNENI, as an example.
+Then ProteinDT adopts the prompts with keywords “more $\alpha$-helices” and “more $\beta$-sheets”, the edited protein structures (edited sequences folded with ESMFold) are shown in Figure 3 (b) and (e).
+Similarly for Figure 3 (c) and (f) when using prompts like “less $\alpha$-helices” and “less $\beta$-sheets”.
+For peptide editing illustration, we take PDB 3IQI [47], Haemophilus influenzae O-acetylserine sulfhydrylase in complex with the peptide MNENI, as an example.
 The input peptide is shown in Figure 3 (g), and the output peptides with higher and lower binding affinity are shown in Figure 3 (h, i).
@@
 Table 2: The satisfactory hit ratio (%) of text-guided protein editing tasks. The best results and the second-best results are marked.
-Structure: $\alpha$-helices Structure: $\beta$-sheets Stability: Villin Stability: Pin1 binding editing more less more less more stable less stable more stable less stable higher lower BERT DSSP BERT DSSP BERT DSSP BERT DSSP BERT pLDDT BERT pLDDT BERT pLDDT BERT pLDDT BindingNet GNINA BindingNet GNINA Galactica 5.46 7.02 13.26 10.53 1.95 2.14 11.50 11.31 7.81 7.81 73.44 67.19 1.41 2.82 8.45 7.04 26.61 21.77 33.87 25.00 ChatGPT 5.46 6.82 7.80 6.43 4.68 2.92 5.07 3.51 29.69 17.19 78.12 18.75 28.17 18.31 39.44 36.62 16.94 12.90 26.61 35.48 ProteinDT (Ours) Latent Interpolation ProteinDiff-RNN 67.64 75.63 19.49 10.14 27.88 11.31 53.61 69.98 17.19 1.56 82.81 98.44 4.23 4.23 95.77 94.37 58.87 29.84 37.90 70.16 ProteinDiff-BERT 38.79 51.27 48.15 34.70 39.38 15.79 42.30 65.50 9.38 1.56 93.75 98.44 2.82 5.63 97.18 94.37 50.81 29.03 43.55 67.74 AR-BERT 27.49 45.61 56.92 39.38 9.16 5.26 69.79 73.10 6.25 1.56 93.75 98.44 1.41 7.04 98.59 92.96 53.23 29.84 45.97 71.77 Latent Optimization 51.66 54.19 33.92 32.94 46.59 39.57 39.38 46.59 50.00 29.69 54.69 76.56 52.11 42.25 47.89 56.34 39.52 37.90 60.48 61.29
+Structure: $\alpha$-helices Structure: $\beta$-sheets Region more less more less ordered disordered BERT DSSP BERT DSSP BERT DSSP BERT DSSP pLDDT pLDDT Galactica 5.46 7.02 13.26 10.53 1.95 2.14 11.50 11.31 8.47 23.73 ChatGPT 5.46 6.82 7.80 6.43 4.68 2.92 5.07 3.51 6.78 62.71 ProteinDT (Ours) Latent Interpolation ProteinDiff-RNN 67.64 75.63 19.49 10.14 27.88 11.31 53.61 69.98 5.08 94.92 ProteinDiff-BERT 38.79 51.27 48.15 34.70 39.38 15.79 42.30 65.50 3.39 94.92 AR-BERT 27.49 45.61 56.92 39.38 9.16 5.26 69.79 73.10 15.25 84.75 Latent Optimization 51.66 54.19 33.92 32.94 46.59 39.57 39.38 46.59 6.78 91.53 Stability: Villin Stability: Pin1 binding editing more stable less stable more stable less stable higher lower BERT Rosetta BERT Rosetta BERT Rosetta BERT Rosetta BindingNet GNINA BindingNet GNINA Galactica 7.81 51.56 73.44 31.25 1.41 4.23 8.45 4.23 26.61 21.77 33.87 25.00 ChatGPT 29.69 57.81 78.12 28.12 28.17 19.72 39.44 71.83 16.94 12.90 26.61 35.48 ProteinDT (Ours) Latent Interpolation ProteinDiff-RNN 17.19 50.00 82.81 54.69 4.23 26.76 95.77 70.42 58.87 29.84 37.90 70.16 ProteinDiff-BERT 9.38 23.44 93.75 68.75 2.82 16.90 97.18 84.51 50.81 29.03 43.55 67.74 AR-BERT 6.25 40.62 93.75 59.38 1.41 25.35 98.59 73.24 53.23 29.84 45.97 71.77 Latent Optimization 50.00 53.12 54.69 40.62 52.11 35.21 47.89 63.38 39.52 37.90 60.48 61.29
 Figure 3: Visual analysis on text-guided protein editing with latent optimization.
@@
 We consider the following six tasks gathered from TAPE [48].
-Secondary structure (SS) prediction [43] is a token-level prediction task, and the goal is to predict the secondary structure (helix, strand, etc.) for each amino acid.
+Secondary structure (SS) prediction [49] is a token-level prediction task, and the goal is to predict the secondary structure (helix, sheet, etc.) for each amino acid.
 SS-Q3 and SS-Q8 refer to the three-way and eight-way classifications, respectively.
@@
 Stability [40] is to predict a proxy for protein stability.
-Both Fluorescence and Stability are regression tasks, and they are measured using Spearman’s $\rho$ metric.
-More details of the datasets are in Supplementary B.
+Both tasks are measured using Spearman’s $\rho$ metric.
 Baselines.
@@
 The property prediction results are in Table 3.
-We show two ProtBERT models followed with ProteinCLAP pretraining, corresponding to the EBM-NCE loss and InfoNCE loss, respectively.
+We show two variants of ProteinCLAP pretraining, corresponding to the EBM-NCE loss and InfoNCE loss, respectively.
 We can observe that both ProteinCLAP pretrained representations can consistently reach the best performance on three structure-related prediction tasks.
-Yet for the other two tasks, Homology and Stability, they are still within the top-three models.
+Yet for the other two tasks, Homology and Stability, they are still within the top three models.
 The only exception is the Fluorescence task, where all the ProtBERT models are performing worse than other baselines.
-Noticeably, ProteinCLAP outperforms OntoProtein on five out of six tasks.
 Table 3: Results on six protein property prediction tasks from TAPE.
@@
 Discussion
-We acknowledge two key limitations that may hinder ProteinDT for broader usage, as discussed below.
 Challenge regarding dataset quality.
@@
 It has 441K data pairs, yet such a dataset size is small compared to the vision and language domains.
-If we want to take the protein backbone structure into consideration, that would be reduced to merely 45K pairs of data [100].
+If we want to take the experimental protein backbone structure into consideration, that would be reduced to merely 45K pairs of data [58].
 Thus, data insufficiency has become the bottleneck of this research direction.
@@
 Challenge regarding evaluation.
-For the text-guided protein design tasks, the textual description is quite flexible, but the output proteins can be tested.
+For the text-guided protein design tasks, the textual description is quite flexible, but the output proteins need to be tested.
 Although most protein properties can be tested experimentally, the cost and throughput of the experiments can vary considerably depending on the property.
-Generative models in the art, language, and vision domains are often subjective and can be evaluated by a human reviewer.
-In contrast, protein engineering is objective but requires wet lab testing for the most rigorous model assessment.
 In this work, we take the computationally feasible solutions with regressor models as surrogate oracles, following the evaluation approach of reinforcement learning models [60].
-However, the validity of such surrogate models is not guaranteed.
-We attempt to reduce our dependence on surrogate models during the editing evaluation by including alternative evaluations from methods such as ESMFold and GNINA, but those tools still do not provide biological ground truth.
+However, the reliability of such surrogate models is not guaranteed.
+To mitigate this dependency, we incorporate alternative evaluation methods, such as ESMFold and GNINA, during the editing assessment.
 Although the preliminary protein editing results are quite encouraging, additional evaluations will be required to more thoroughly assess the strengths and weaknesses of ProteinDT on diverse editing tasks.
-Conclusion.
-In this work, we introduced a text-guided protein design framework named ProteinDT.
-ProteinDT trains a ProteinCLAP model, a ProteinFacilitator model, and a decoder model (AR and ProteinDiff), to bridge the gap between the natural language and protein sequence structure.
-We examined and developed three downstream tasks to verify the versatile functionalities of ProteinDT: text-to-protein generation, zero-shot text-guided protein editing, and protein property prediction.
-Societal impacts.
-Although text-based protein design has many potential positive applications in agriculture, bioengineering, and therapeutics, it can be considered a dual-use technology.
-Much like generative models for small molecules [56], ProteinDT could be applied to generate toxic or otherwise harmful protein sequences.
-Even though acting on these designs would require a wet lab, synthesizing custom amino acid sequences is typically straightforward compared to synthesizing novel small molecules.
-These potential negative impacts in the current version of ProteinDT may be limited by the text in the SwissProtCLAP dataset because the model is not expected to perform well for terms and concepts not represented in UniProt.
-Our initial manual review of a sample of SwissProtCLAP indicated it does contain phrases describing toxic proteins (e.g., “May have toxic activity.
-Expressed by the venom gland.”) but other potentially harmful terms examined (e.g., “dangerous”) were used in different contexts that do not pose a risk to human health (e.g., “when trehalose levels in the cell become dangerously low, TreS could expedite the conversion of glycogen to maltose via its amylase activity and then convert the maltose to trehalose”).
-Future efforts to expand the training dataset and model improvements could increase the dual-use risk.
 Outlook.
@@
 ProteinDT shows a potential strategy for jump-starting future protein design of specific functions without requiring experimental data by leveraging the rich information in the textual description.
-Method
+Methods
 Backgrounds on Generative Models
@@
 Concretely, the ProteinFacilitator model explicitly produces the protein sequence representation from the text representation, and we use a Gaussian distribution to model such a process.
+Note that this is just a simplifying choice of assumption and not a limitation of our method.
+We have text-to-protein generation and text-guided protein editing results in the Results Section (Table 1-2, Figure 3) verifying that such an assumption can give us reasonably good results.
 We introduce an alignment module $f_{\text{ProteinFacilitator}}(\cdot)$ and then we estimate the ProteinFacilitator distribution using a conditional Gaussian as $p({\bm{z}}_{\text{p}}|{\bm{z}}_{\text{t}})=\mathcal{N}({\bm{z}}_{\text{p}};f_{\text{ProteinFacilitator}}({\bm{z}}_{\text{t}}),I)$.
-To minimize its likelihood, the objective is essentially the representation reconstruction (regression) task [72]:
+To minimize its likelihood, the objective is essentially the representation reconstruction (regression) task [76, 18]:
 $\mathcal{L}_{\text{ProteinFacilitator}}=\|f_{\text{ProteinFacilitator}}({\bm{z}}_{\text{t}})-{\bm{z}}_{\text{p}})\|^{2},$ (6)
@@
 In the Results Section, we empirically verify the effectiveness of introducing the ProteinFacilitator model on the text-to-protein design task.
-Protein Decoding
+Protein Decoder Model
 With the condition ${\bm{z}}_{\text{c}}$ generated by the aforementioned ProteinFacilitator, a decoder model is applied to conduct conditional generative modeling for protein generation, i.e., $p({\bm{x}}_{\text{p}}|{\bm{z}}_{\text{c}})$.
@@
 We repeat such sampling process for $T$ timesteps, from $\tilde{\bm{x}}_{T}$ to $\tilde{\bm{x}}_{0}$, which yields protein sequence generation.
+Code Availability
+The source code can be found at this GitHub repository or Zenodo [88].
+Data Availability
+The dataset generation scripts can be found at this GitHub repository. We also provide the preprocessed pretraining dataset (SwissProtCLAP) at this HuggingFace link.
 Acknowledgements
@@
 program at Mila-UdeM, and was supported in part by the Natural Sciences and Engineering Research Council (NSERC) Discovery Grant, the Canada CIFAR AI Chair Program, collaboration grants between Microsoft Research and Mila, Samsung Electronics Co., Ltd., Amazon Faculty Research Award, Tencent AI Lab Rhino-Bird Gift Fund, two NRC Collaborative R&D Projects, IVADO Fundamental Research Project grant PRF-2019-3583139727, and NSF award CHE 2226451.
-Author Contributions Statement
-S.L., Y.L., A.G., Y.Z., Z.X., W.N., A.R., C.X., J.T., H.G., and A.A.
-conceived and designed the experiments.
-S.L.
-performed the experiments.
-S.L., Y.L., and A.G.
-analyzed the data.
-S.L., Y.L., Z.L., and J.L.
-contributed analysis tools.
-S.L., Y.L., Z.L., A.G., C.X., H.G., and A.A.
-wrote the paper.
-C.X., J.T., H.G., and A.A.
-contributed equally to advising this project.
-Competing Interests Statement
-The authors declare no competing interests.
-Supplementary Information
-Appendix A Related Work
-A.1 Multi-modality Learning
-Recently, there has been a resurgent research interest in multi-modal learning in the machine learning community, exemplified by CLIP [86].
-CLIP bridges the gap between image and language modalities by aligning the representation of the two via contrastive learning.
-This operation opens a novel research track on text-to-image (controllable) generation with natural language [87, 88, 89, 90, 91, 92, 93].
-Multi-modality learning on small molecules.
-One active research line is utilizing external information to help augment the molecule representation learning, e.g., using a biological knowledge graph [94] or biomedical textual information [95, 96, 97] to augment the molecule representation learning.
-SGNN-EBM [94] and Text2Mol [95] focus on the supervised setting, while KV-PLM [96] aims at pre-training with a masked auto-encoding paradigm.
-MoMu [98] follows the setting of KV-PLM and further tests the text-to-molecule generation as in Text2Mol.
-MoleculeSTM [97] applies the contrastive learning paradigm to align the two modalities and conducts the text-guided molecule editing task.
-It reveals the superior understandability of natural language for finding novel biochemical concepts.
-Multi-modality learning on macromolecules.
-For macromolecules like proteins, OntoProtein [99] first combines the protein sequence and textual description into a biological knowledge graph (KG), and the pretraining is done by the relation prediction on this biological KG.
-Chroma [100] is a parallel work that also conducts text-guided protein programmable editing, the ProCap model.
-Meanwhile, there are fundamental differences between ProCap and ProteinDT, as discussed in Table 4.
-A.2 Protein Sequence and Structure Design
-Generative protein sequence models and their applications were recently reviewed by [101].
-Many recent protein sequence design models generate the protein sequence in an autoregressive manner [102, 103, 104, 105], i.e., amino-acid by amino-acid.
-Other protein sequence generation models incorporate autoencoders [106, 107, 108, 109, 110], generative adversarial networks [111, 112], or iterative masking with multiple sequence alignments [113].
-Alternatively, reinforcement learning can also be used to generate sequences [114, 115, 116] or optimize sequences [117, 118, 119].
-The reinforcement learning optimization setting resembles the ProteinDT text-guided protein editing task in that they have the same goal of improving an initial sequence.
-However, the reinforcement learning approach usually requires labels from experimental data unless the target property can be simulated [117].
-In contrast, ProteinDT is a zero-shot approach requiring no experimental labels, only a suitable text prompt.
-Other generative models include structure in the design process [120].
-One approach is to start from a 3D protein backbone structure and design a compatible amino acid sequence [121, 122, 123, 124], which is known as inverse folding.
-Alternatively, there are also diverse techniques for designing 3D structures [125, 126, 127, 100, 128, 129, 130].
-In Table 4, we provide a specific comparison between ProCap (Chroma) and ProteinDT.
-Table 4: Comparison between ProCap (Chroma) and ProteinDT.
-ProCap (Chroma) ProteinDT Data Structure ProCap or Chroma is focusing on the protein backbone structure.
-ProteinDT is focusing on the protein sequence.
-Dataset 45K text-backbone pairs from PDB and UniProt.
-441K text-sequence pairs from SwissProtCLAP.
-Modeling ProCap builds a protein-to-text model (protein captioning), $p({\bm{x}}_{\text{t}}|{\bm{x}}_{\text{p},3D})$, where ${\bm{x}}_{\text{t}}$ is the textual description and ${\bm{x}}_{\text{p},3D}$ is the backbone (3D) structure.
-ProteinDT is mainly focusing on the text-to-protein model, $p({\bm{x}}_{\text{p}}|{\bm{x}}_{\text{t}})$, where ${\bm{x}}_{\text{p}}$ is the protein sequence.
-Inference ProCap utilizes classifier guidance [131], which is the MCMC sampling in a Bayes manner.
-The inference in ProteinDT depends on the decoder model.
-In the current version, autoregressive sampling and MCMC sampling are considered.
-Framework ProCap is specifically designed for a denoising diffusion generative model.
-ProteinDT is model-agnostic and thus is more flexible.
-Other deep generative models (energy-based model, variational auto-encoder, flow-based model, etc) can also be adopted.
-Appendix B Pretraining Dataset SwissProtCLAP
-We list the key statistics of the SwissProtCLAP dataset (as of date 2022.09.08) in Table 5.
-There are many redundant gene names, and we emphasize this by explicitly listing the number of protein identifiers and the number of unique gene names.
-The organism count is based on the number of unique NCBI Taxonomy identifiers.
-Table 5: Statistics on SwissProtCLAP: the # text-protein pairs, # protein identifiers, # unique gene names, and # unique organisms.
-# text-protein pairs # proteins # genes # organisms 441,013 441,013 327,577 13,339
-Then we randomly select six text-protein pairs from SwissProtCLAP in Table 6 for illustration, so as to give a better intuition about the dataset contents.
-Table 6: Examples of text-protein pairs in SwissProtCLAP.
-Protein Sequence Text Sequence MSLNAEQKAKVVLEHGSSAHDTGSTEVQVALL
-TLRINDLQKHFLEHKKDHHSRRGLLRMVSQRR
-KLLDYLKKRNISKYTDLIQSLGLRK
-One of the primary rRNA binding proteins, it binds directly to 16S rRNA where it helps nucleate assembly of the platform of the 30S subunit by binding and bridging several RNA helices of the 16S rRNA.
-Forms an intersubunit bridge (bridge B4) with the 23S rRNA of the 50S subunit in the ribosome.
-Part of the 30S ribosomal subunit.
-Forms a bridge to the 50S subunit in the 70S ribosome, contacting the 23S rRNA.
-Belongs to the universal ribosomal protein uS15 family.
-MRATVGLVEAIGIRELRQHASRYLARVEAGEE
-LGVTNKGRLVARLIPVQAAERSREALIESGVL
-IPARRPQNLLDVTAEPARGRKRTLSDVLNEMR
-Belongs to the phD/YefM antitoxin family.
-LKPDEELQGPGGVLSRGYFVFRPRN
-Stimulates uterine smooth muscle contraction and causes selective vasoconstriction. Belongs to the NmU family.
-MARSLKKGPFVDHHLAKKVESAAGSKKPIKTW
-SRRSMILPEMVGITIAVHNGKNHIPVLVNENM
-VGHKLGEFAVTRTFKGHGGDKKSSR
-Protein S19 forms a complex with S13 that binds strongly to the 16S ribosomal RNA. Belongs to the universal ribosomal protein uS19 family.
-MAKVIIEIKNTVSGIKGRNLRTSIAVDGSAEL
-DGDEGTLAGMVALLVLNKSQKIINESAHEAIE
-ILKNDGVITSGRVTEMAVEKTCH
-Expressed in the early phase of the viral replicative cycle. Expression of early genes is repressed by viral Repc (latency) and favored by viral Ner protein.
-MSRTIFCTFLNKEADGLDFQLYPGELGKRIFN
-EISKEAWGQWMAKQTMLINEKKLNTMNPDDRK
-LLEQEMVRFLFEGHDVHIDGYTPPEK
-Could be a mediator in iron transactions between iron acquisition and iron-requiring processes, such as synthesis and/or repair of Fe-S clusters in biosynthetic enzymes.
-Monomer.
-Belongs to the Fe(2+)-trafficking protein family.
-Appendix C Architecture Details and Ablation Studies
-We list the implementation details of ProteinDT, following the pretraining steps and downstream tasks.
-C.1 ProteinCLAP Model
-We use two BERT models to model the protein sequence and text sequence, respectively.
-Then we take the SciBERT [132] and the ProtBERT [102] as the starting checkpoint, followed by ProteinCLAP pretraining.
-We also add a prediction head to each of the BERT models since this approach has been found [133, 134] empirically useful in the contrastive learning setting.
-The key hyperparameters are listed below.
-Table 7: Hyperparameter specifications for ProteinCLAP.
-Hyperparameter Value epochs {5, 10} learning rate for text sequence representation {1e-5, 1e-6} learning rate for protein sequence representation {1e-5, 1e-6} latent dimension of prediction head {256} objective function {EBM-NCE, InfoNCE} temperature {0, 0.1 } batch size {8}
-C.2 ProteinFacilitator Model
-We consider using the Gaussian distribution as the ProteinFacilitator distribution in ProteinDT. The key hyperparameters are listed below.
-Table 8: Hyperparameter specifications for ProteinFacilitator distribution.
-Hyperparameter Value epochs {32} learning rate {1e-6} objective function {MSE} batch size {16}
-C.3 Decoder Model
-(a) Autoregressive model.
-(b) Diffusion model.
-Figure 4: The inference illustration for two decoder models (step 3 in ProteinDT).
-Figure 4(a) is an autoregressive (AR) model based on the Transformer.
-It generates the protein sequence token-by-token.
-Figure 4(b) is ProteinDiff, and a Transformer-based transition network is displayed.
-It first randomly samples a noised sequence, then conducts a denoising process using the transition network.
-For the decoder, we consider two categories of models: the autoregressive (AR) and the discrete diffusion model (ProteinDiff).
-We list the pipeline in Figure 4, and the key hyperparameters are listed below.
-Table 9: Hyperparameter specifications for decoder distribution.
-model Hyperparameter Value Autoregressive epochs {10} learning rate {1e-4, 1e-5} ProteinDiff transition network architecture {RNN, Transformer} epochs {10} learning rate {1e-4, 1e-5} hidden dimension {16, 32}
-Inference for AR
-For the autoregressive (AR) model, we consider two variants in inference: the greedy search and the beam search.
-Inference for ProteinDiff
-The reverse process $\tilde{\bm{x}}_{T}\to\tilde{\bm{x}}_{0}$ in the diffusion model corresponds to the protein sequence sampling for protein generation in ProteinDiff.
-In detail, to generate a protein sequence, we first generate a diffused sequence ${\bm{x}}_{T}$ with all tokens being the “[MASK]” token.
-Then we consider two sampling strategies, as below.
-(1) The simplified sampling [135]
-For timestep $t$, we sample the mask positions with probability $1/t$ and predict the masked tokens following the distribution $p_{\theta}(\tilde{\bm{x}}_{0}|\tilde{\bm{x}}_{t+1},{\bm{z}}_{\text{c}})$.
-This is an approximation sampling, yet works quite well on discrete data, like quantized codes.
-(2) The weighted sampling [136]
-For timestep $t$, the $p_{\theta}(\tilde{\bm{x}}_{t}|\tilde{\bm{x}}_{t+1},{\bm{z}}_{\text{c}})$ is indeed $p_{\theta}(\tilde{\bm{x}}_{0}|\tilde{\bm{x}}_{t+1},{\bm{z}}_{\text{c}})$ weighted by a posterior distribution.
-More concretely, first, with the predefined transition matrix and forward process function, we can get the posterior $q(\tilde{\bm{x}}_{t}|\tilde{\bm{x}}_{t+1},\tilde{\bm{x}}_{0})$ in a closed-form as:
-$q(\tilde{\bm{x}}_{t}|\tilde{\bm{x}}_{t+1},\tilde{\bm{x}}_{0})=\frac{q(\tilde{\bm{x}}_{t+1}|\tilde{\bm{x}}_{t})\cdot q(\tilde{\bm{x}}_{t}|\tilde{\bm{x}}_{0})}{q(\tilde{\bm{x}}_{t+1}|\tilde{\bm{x}}_{0})}=\text{Categorical}\Big{(}\tilde{\bm{x}}_{t};p=\frac{\tilde{\bm{x}}_{t+1}Q_{t+1}^{T}\odot\tilde{\bm{x}}_{0}\bar{Q}_{t}}{\tilde{\bm{x}}_{0}\bar{Q}_{t+1}\tilde{\bm{x}}_{t+1}^{T}}\Big{)},$ (9)
-where $\bar{Q}_{t}=\prod_{s=1}^{t}Q_{s}$, and $\odot$ is the element-wise product.
-Then we can get the logits $\tilde{\bm{x}}_{0}$ using the transition network (same as the simplified sampling), i.e., $p_{\theta}(\tilde{\bm{x}}_{0}|\tilde{\bm{x}}_{t},{\bm{z}}_{\text{c}})$.
-Thus the weighted sampling for $\tilde{\bm{x}}_{t}$ is:
-$p_{\theta}(\tilde{\bm{x}}_{t}|\tilde{\bm{x}}_{t+1},{\bm{z}}_{\text{c}})=\sum_{\bar{\bm{x}}_{0}}q(\tilde{\bm{x}}_{t}|\tilde{\bm{x}}_{t+1},\bar{\bm{x}}_{0})\cdot p_{\theta}(\bar{\bm{x}}_{0}|\tilde{\bm{x}}_{t+1},{\bm{z}}_{\text{c}}),$ (10)
-where $\bar{\bm{x}}_{0}$ is the one-hot representation of $\tilde{\bm{x}}_{0}$.
-The effect of sampling strategies.
-Our ablation studies, detailed in Table 10, indicate that there isn’t a significant disparity between the two sampling methods.
-Notably, employing RNN with a simplified sampling method and utilizing the backbone yields preferable outcomes, whereas Transformer-based ProteinDiff performs better with weighted sampling.
-Hence, for the sake of simplicity, we present the results exclusively using simplified sampling.
-This finding holds promise for future explorations in this domain.
-Table 10: Ablation studies on using simplified and weighted sampling in ProteinDiff for text-to-protein-generation. The retrieval accuracies (%) are reported with different $T$-choose-1 tasks.
-Model $T=4$ $T=10$ $T=20$ ProteinDiff-RNN simplified sampling 40.50 21.50 15.00 weighted sampling 33.50 17.00 13.00 ProteinDiff-Transformer simplified sampling 51.50 25.00 13.50 weighted sampling 53.50 28.50 21.00
-Appendix D Task Specification
-D.1 Downstream Task: Text-to-Protein Generation
-As discussed in the main manuscript, one of the most important tricks is whether to not to use ProteinFacilitator.
-In this section, we would like to discuss other important tricks and hyperparameters for text-to-protein generation.
-D.1.1 Sequence Length
-For the AR model, there is an end-of-sequence (“EOS”) token explicitly telling the end of generated sequences.
-For ProteinDiff, we either take the first padding token (“PAD”) as the end of the protein sequence, or we ignore all the padding tokens in the generated protein sequences.
-D.1.2 Selection of Generated Protein Sequences
-Note that we also incorporate ProteinCLAP to help augment the quality of generated protein sequences.
-Concretely, for each textual description, we generate 16 protein sequences, then we select the one with the highest similarity with the textual description.
-We conduct this for all the baselines and ProteinDT.
-D.2 Downstream Task: Text-guided Protein Editing
-Zero-shot text-guided protein editing is one of the most important tasks we design. Due to the space limitation in the main manuscript, we would like to provide more details in this section.
-D.2.1 Editing Models
-We cover two editing methods: latent interpolation and latent optimization.
-Latent interpolation does not require extra training, yet latent optimization requires us to train a decoder $g$.
-We list the detailed comparison between the two in Table 11.
-Specifically in latent optimization, we consider a recurrent neural network (RNN) model, and the hyperparameters are listed in Table 12.
-Determining the appropriate decoder in latent optimization remains an unresolved query.
-Within ProteinDT, our initial step in this avenue of research involves the utilization of RNN, recognized as a highly efficient and proficient sequential model.
-Empirical evidence substantiates the commendable performance of RNN as a decoder in executing protein design tasks.
-It is imperative to note that the investigation into more sophisticated decoder architectures is deferred for future exploration.
-Table 11: Comparison of key modules between latent interpolation and latent optimization.
-Latent Interpolation Latent Optimization Representation Sequence-level Token-level Decoding Token-by-token Token Prediction Learning No Learning Learning Required Text-guidance SLERP Interpolation Similarity
-Table 12: Hyperparameter specifications for latent optimization decoder (RNN).
-Hyperparameter Value epochs {10} learning rate {1e-3, 5e-3} RNN layer {2, 3}
-D.2.2 Dataset of Editing Tasks
-In this section, we would like to provide more details on the dataset of each editing task. All the data preprocessing scripts can be found in the GitHub repository.
-Structure editing.
-We follow prior work on protein secondary structure prediction [137].
-The oracle evaluator (as will be discussed next) is pretrained on the training set, so we take the test set as input protein sequences (data splitting from the original paper).
-Three distinct test sets are available, and in accordance with [138], we opt for the “cb513” subset, comprising 513 protein sequences.
-This holds for the editing tasks on both alpha-helix and beta-sheet.
-Stability editing.
-Rocklin et al.
-assessed protein stability with a high-throughput assay based on protease resistance [137].
-Cells were treated with multiple concentrations of a protease.
-A protein’s measured ability to tolerate the protease was normalized by its estimated tolerance in an unfolded state to generate a stabilty score.
-Therefore, larger positive stability scores represent more stable proteins.
-Negative scores mean that the unfolded state was estimated to be more resistant to the protease treatment than what was measured experimentally for that protein.
-Although Rocklin et al.
-and the TAPE [139] stability dataset derived from this study contain stability measurements for multiple rounds of iterative protein design as well as saturation mutagenesis experiments for designed and natural proteins, we focus on two natural proteins, villin HP35 and Pin1 WW domain.
-In this controlled setting, we increase the likelihood that the regressor will be able to estimate reasonable stability scores for the output edited sequences.
-If we were to instead train the regressor on all available sequences and edit a wide variety of input sequences, we may be less confident in the regressor’s ability to score those generated sequences accurately.
-We extract the villin saturation mutagenesis stability data from the TAPE dataset, which places all of these sequences in the test split, by selecting all 631 instances with the topology set to villin.
-The saturation mutagenesis data contains the wild type villin HP35 sequence and all 18 possible single amino acid mutations at each position in the sequence, excluding cysteine (C).
-We remove the five-character prefix GSSGS and three-character suffix GSS from the protein sequences before training the regressor and providing the sequeces as input to the editing model.
-The prefix and suffix represent padding residues added during library design [137] and are not part of the 35 amino acids long villin sequence.
-Similarly, we extract Pin1 stability data from the TAPE dataset.
-We select all 703 instances with the topology set to Pin1.
-We remove the one-character prefix G and three-character suffix GSS, leading to 39 amino acids in the Pin1 sequences.
-We do not remove prefix or suffix sequences for the TAPE stability prediction task above.
-Peptide binding editing.
-We follow recent work of estimating the protein-ligand binding affinity and dynamics [140].
-Here we extract the protein complexes with a peptide as the ligand.
-On the other hand, we also need the textual description of the protein sequences.
-To achieve this, we keep the protein-ligand complexes wherein the proteins exist with the SwissProtCLAP.
-Following the filtration process, this results in a total of 124 peptide sequences.
-D.2.3 Sequence Length
-For the length of the generated sequences, we maintain consistency with the strategy used for text-to-protein generation.
-However, an additional constraint is introduced: the maximum sequence length must fall within a specific range relative to the input protein sequences.
-To do so, we dynamically establish the maximum sequence length for generated proteins by adopting the maximum protein sequence length from each batch of protein sequences.
-Namely, we take the maximum protein sequence length for each input batch as the upper limit for the generated protein sequences.
-Furthermore, in the context of the stability editing and peptide binding editing tasks, the input protein sequences are relatively compact, reaching a maximum of 39 and 15 amino acids, respectively.
-However, for the structure editing task, the computational burden incurred during the folding evaluation is considerable.
-Consequently, we take all the protein sequences with length over 512 as invalid proteins.
-D.2.4 Selection of Generated Protein Sequences
-Additionally, we incorporate ProteinCLAP to enhance the quality of generated protein sequences.
-In alignment with text-to-protein generation, we generate 16 protein sequences for each input protein sequence and text prompt.
-Subsequently, we select the sequence exhibiting the highest similarity with either the input description or input protein sequence.
-This selection methodology is applied within ProteinDT.
-However, it differs from the baselines (ChatGPT and Galactica), as they amalgamate the input protein sequence and text prompt into a singular sentence, and such a design framework does not support the post-sampling selection process outlined above.
-It’s crucial to note that the primary distinction between editing and text-to-protein generation lies in the uniformity of inputs for the latter across all baselines and ProteinDT.
-D.2.5 Evaluation Function
-Next, we would like to list more details on the oracle evaluation function of different editing tasks.
-Structure editing.
-One oracle evaluation function we considered is a protein sequence evaluator for secondary structure prediction (as discussed in the property prediction downstream below).
-More concretely, we are using the ProteinDT-ProteinCLAP-EBM-NCE, which is the best predictor of SS-Q3 in our evaluations.
-Stability editing.
-The oracle evaluation function we considered is the optimal protein sequence evaluator for stability (as discussed in the property prediction downstream below).
-More concretely, we are using ProtBERT, which is the best predictor of Stability in our evaluations.
-Peptide binding editing.
-For the peptide editing, we first train an oracle model on the protein-peptide binding datasets from PDB.
-The geometric model satisfies the SE(3)-equivariance using the vector frame basis [141], which is one of the state-of-the-art models on protein structure modeling.
-Then we adopt GNINA as one of the oracle evaluation metrics [142], which is a docking software that uses an ensemble of convolutional neural networks as the scoring function.
-For each conformation, it simultaneously predicts the pose score (probability that the current pose resembles the actual pose) and binding affinity.
-An ideal docking conformation should have a high affinity under the pose score of 1.
-Thus, the final docking score is calculated as the product of the pose score and binding affinity, as the evaluation of the current docking conformation quality.
-Satisfactory function.
-The oracle evaluation functions listed above are for evaluating one protein sequence (either input protein sequence or output protein sequence).
-The next step is to calculate the satisfactory hit ratio (%), which measures the ratio of the output score being higher or lower than the input score, as indicated in the text prompt.
-D.2.6 Text Prompt in Text-guided Protein Editing
-In this section, we add details of the text prompts for 10 editing tasks with ProteinDT and two baselines.
-Table 13: Prompts for structure editing. The curly brackets “{}” represent the input protein sequence.
-Task Method Prompt more alpha helices Galactica Given an input amino acid sequence [START_AMINO]{}[END_AMINO].
-Can you modify the amino acid sequence to have more alpha helices in the secondary, which needs to be similar to the input sequence?
-[START_AMINO] ChatGPT Given an input amino acid sequence {}.
-Can you modify the amino acid sequence to have more alpha helices in the secondary structure, which needs to be similar to the input sequence (just AA sequence, and no explanation)?
-ProteinDT modify the amino acid sequence to have more alpha helices in the secondary structure less alpha helices Galactica Given an input amino acid sequence [START_AMINO]{}[END_AMINO].
-Can you modify the amino acid sequence to have less alpha helices in the secondary, which needs to be similar to the input sequence?
-[START_AMINO] ChatGPT Given an input amino acid sequence {}.
-Can you modify the amino acid sequence to have less alpha helices in the secondary structure, which needs to be similar to the input sequence (just AA sequence, and no explanation)?
-ProteinDT modify the amino acid sequence to have less alpha helices in the secondary structure more beta sheets Galactica Given an input amino acid sequence [START_AMINO]{}[END_AMINO].
-Can you modify the amino acid sequence to have more beta sheets in the secondary, which needs to be similar to the input sequence?
-[START_AMINO] ChatGPT Given an input amino acid sequence {}.
-Can you modify the amino acid sequence to have more beta sheets in the secondary structure, which needs to be similar to the input sequence (just AA sequence, and no explanation)?
-ProteinDT modify the amino acid sequence to have more beta sheets in the secondary structure less beta sheets Galactica Given an input amino acid sequence [START_AMINO]{}[END_AMINO].
-Can you modify the amino acid sequence to have less beta sheets in the secondary, which needs to be similar to the input sequence?
-[START_AMINO] ChatGPT Given an input amino acid sequence {}.
-Can you modify the amino acid sequence to have less beta sheets in the secondary structure, which needs to be similar to the input sequence (just AA sequence, and no explanation)?
-ProteinDT modify the amino acid sequence to have less beta sheets in the secondary structure
-Table 14: Prompts for stability editing. The following are used for both villin and Pin1, and the curly brackets “{}” represent the input protein sequence.
-Task Method Prompt higher stability Galactica Given an input amino acid sequence [START_AMINO]{}[END_AMINO].
-Can you modify the amino acid sequence to have higher stability, which needs to be similar to the input sequence?
-[START_AMINO] ChatGPT Given an input amino acid sequence {}.
-Can you modify the amino acid sequence to have higher stability, which needs to be similar to the input sequence (just AA sequence, and no explanation)?
-ProteinDT modify the amino acid sequence to have higher stability lower stability Galactica Given an input amino acid sequence [START_AMINO]{}[END_AMINO].
-Can you modify the amino acid sequence to have lower stability, which needs to be similar to the input sequence?
-[START_AMINO] ChatGPT Given an input amino acid sequence {}.
-Can you modify the amino acid sequence to have lower stability, which needs to be similar to the input sequence (just AA sequence, and no explanation)?
-ProteinDT modify the amino acid sequence to have lower stability
-Table 15: Prompts for peptide binding editing.
-The curly brackets “{}” represent each binding task’s input peptide sequence and protein property description accordingly.
-The detailed descriptions of each protein target can be found in the GitHub repository.
-Task Method Prompt higher affinity Galactica Given an input peptide amino acid sequence [START_AMINO]{}[END_AMINO] and a target protein.
-The target protein satisfies the following property.
-{} Can you modify the peptide amino acid sequence to have higher binding affinity with the target protein, which needs to be similar to the input sequence?
-[START_AMINO] ChatGPT Given an input peptide amino acid sequence {} and a target protein.
-The target protein satisfies the following property.
-{} Can you modify the peptide amino acid sequence to have higher binding affinity with the target protein, which needs to be similar to the input sequence (just AA sequence, and no explanation)?
-ProteinDT modify the peptide amino acid sequence to have higher binding affinity with the target protein.
-The target protein satisfies the following property.
-{} lower affinity Galactica Given an input peptide amino acid sequence [START_AMINO]{}[END_AMINO] and a target protein.
-The target protein satisfies the following property.
-{} Can you modify the peptide amino acid sequence to have lower binding affinity with the target protein, which needs to be similar to the input sequence?
-[START_AMINO] ChatGPT Given an input peptide amino acid sequence {} and a target protein.
-The target protein satisfies the following property.
-{} Can you modify the peptide amino acid sequence to have lower binding affinity with the target protein, which needs to be similar to the input sequence (just AA sequence, and no explanation)?
-ProteinDT modify the peptide amino acid sequence to have lower binding affinity with the target protein.
-The target protein satisfies the following property.
-{}
-D.3 Downstream Task: Property Prediction
-The detailed specifications for the six TAPE [139] tasks can be found in Table 16.
-Table 16: Dataset specification.
-dataset type # train # val #test # labels
-SS-Q3 token-level classification 8,678 2,170 649 3
-SS-Q8 token-level classification 8,678 2,170 649 8
-Contact token-level (token pair) classification 24,651 224 40 2
-Homology sequence-level classification 12,312 736 1,976 1,195
-Fluorescence sequence-level regression 21,446 5,362 27,217 -
-Stability sequence-level regression 53,614 2,512 12,851 -
-We follow the same hyperparameters in OntoProtein [99], as shown in Table 17.
-Table 17: Hyperparameter specifications for protein property prediction.
-Task epoch batch size warmup ratio learning rate frozen BERT Optimizer SS-Q3 5 32 0.08 3e-5 False AdamW SS-Q8 5 32 0.08 3e-5 False AdamW Contact 10 8 0.08 3e-5 False AdamW Homology 10 64 0.08 3e-5 False AdamW Fluorescence 25 64 0 3e-5 True AdamW Stability 5 32 0.08 3e-5 False AdamW